### v1

# To Rely or Not to Rely? Evaluating Interventions for Appropriate Reliance on Large Language Models

Abstract.
As Large Language Models become integral to decision-making, optimism about their power is tempered with concern over their errors. Users may over-rely on LLM advice that is confidently stated but wrong, or under-rely due to mistrust. Reliance interventions have been developed to help users of LLMs, but they lack rigorous evaluation for appropriate reliance. We benchmark the performance of three relevant interventions by conducting a randomized online experiment with 400 participants attempting two challenging tasks: LSAT logical reasoning and image-based numerical estimation. For each question, participants first answered independently, then received LLM advice modified by one of three reliance interventions and answered the question again. Our findings indicate that while interventions reduce over-reliance, they generally fail to improve appropriate reliance. Furthermore, people became more confident after making incorrect reliance decisions in certain contexts, demonstrating poor calibration. Based on our findings, we discuss implications for designing effective reliance interventions in human-LLM collaboration.
1. Introduction
As people increasingly incorporate Large Language Models (LLMs) into their decision-making workflows, there is cause for both tremendous optimism and skeptical concern (Zhao et al., 2023; Eigner and Händler, 2024). LLMs encode remarkably vast amounts of knowledge; can communicate with users in clear, natural language; and have the potential to improve both the quality and efficiency of users’ decision-making (Tankelevitch et al., 2024). However, people can over-rely on LLM advice (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) or, in contrast, under-rely on it if they mistrust their LLM partner (Zhang, 2023). LLMs are prone to hallucination and error, yet often present their outputs with impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023). In contrast, algorithm aversion—the tendency to distrust algorithmic agents—can lead users to ignore helpful recommendations due to general distrust towards the technology (Dietvorst et al., 2015; Burton et al., 2020).
Conscious effort must therefore be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024). Neither over-reliance nor under-reliance is ideal; instead, the focus should be to achieve appropriate reliance (Schemmer et al., 2022; Ma et al., 2023; He et al., 2023). Research in the tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021). In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in human-LLM collaboration.
Due to the risk of hallucinations, reliance and calibration interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024). Reliance interventions that involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024), are being actively developed and evaluated. Since interventions are evolving rapidly, many of these methods still lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022). Furthermore, people’s subjective perceptions—such as their self-confidence—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration. Finally, the impact of different intervention types on reliance and perceptions remains unclear, as existing studies do not benchmark multiple interventions. We seek to address these gaps by asking the following research questions:
RQ1: How do different reliance interventions affect users’ tendencies to over- and under-rely on LLM advice?
RQ2: Can reliance interventions appropriately calibrate users’ confidence in LLM advice?
Summary of Contributions. We conduct a pre-registered, randomized online experiment with participants where they attempted two diverse and challenging question-answering tasks: LSAT logical reasoning and image-based numerical estimation. Participants first attempted the tasks on their own, then were shown LLM advice that was randomly treated with one of four conditions (three reliance interventions and one baseline) and given the question again. To evaluate reliance, we implemented a rigorous set of evaluation metrics that isolate the decision-making process users undergo when choosing to rely on the LLM’s advice or their own judgment.
We find that while interventions can significantly reduce over-reliance, they are generally not effective at improving appropriate reliance. These interventions tend to reduce reliance overall—reducing over-reliance at the expense of useful reliance. Participants were also generally miscalibrated in their confidence, reporting a higher increase in confidence when they make the wrong reliance decision. Out of all the interventions, Reliance Disclaimer is the only one to improve appropriate reliance and better calibrate confidence when participants relied on LLM advice, but this effect was only observed on the LSAT task. Uncertainty Highlighting worsened subjective perceptions and had the least effect on reliance out of the three interventions. Implicit Answer induced better self-reliance in both tasks, but at the cost of requiring significantly more time for participants to process the information. We also discuss the implications of our results for the design and evaluation of appropriate reliance interventions for the future of human-LLM interactions. We highlight the need to perform rigorous, human-centered evaluations of technical LLM calibration methods, particularly with emphasis on achieving appropriate reliance. Furthermore, as the trade-off between balancing over- and under-reliance may be difficult to avoid, defining which outcome to prioritize should be done within the context of the specific human-LLM collaboration task.
2. Related Works
We provide an overview of the growing body of research in human-AI interactions, human-LLM collaboration, and LLM calibration techniques that have been explored in both human-computer interaction (HCI) and AI spaces.
2.1. Appropriate Reliance on AI
The need for appropriate reliance has been a constant over the years of technological development (Lee and See, 2004). With the advancement of AI technologies, much attention has turned on how end users can incorporate assistive AI models into their decision-making tasks (Lai et al., 2021; Kleinberg et al., 2018; Chen et al., 2023). Given that AI predictions are not guaranteed to be accurate, the user must decide whether to or not to incorporate the AI’s advice (Sohrawardi et al., 2024; Chiang and Yin, 2022; Nguyen et al., 2018; Inkpen et al., 2023; Lu and Yin, 2021; Cao and Huang, 2022). The concept of appropriate reliance can typically be defined as ”relying on the AI when it’s correct, and relying on yourself when it’s not”—which concerns reducing both over-reliance and under-reliance (Schemmer et al., 2023; Schoeffer et al., 2024). Root causes for inappropriate reliance can come from misreliance (inherently nonoptimal rate of reliance) or discrimination errors (being unable to identify the payoff of following the AI relative to following oneself) (Guo et al., 2024; Ma et al., 2023).
Factors that influence reliance include the expertise of the user in the task, their cognitive engagement in the task, and the information presented by the AI (Chen et al., 2023; He et al., 2023; Gajos and Mamykina, 2022). On the individual level, people may also be averse or appreciative towards algorithms, perhaps through pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023). A number of strategies have been developed to augment users’ engagement and to provide them with more information of the AI decision, such as through providing explanations (Reingold et al., 2024; Vasconcelos et al., 2023b; Zhang et al., 2020; Schoeffer et al., 2024). The best interventions for appropriate reliance should be designed to support the users’ abilities to verify the AI’s recommendation, not necessarily to provide a fully transparent system (Fok and Weld, 2023; Poursabzi-Sangdeh et al., 2021; Wang and Yin, 2021). We adapt appropriate reliance metrics (Schemmer et al., 2023) to assess the ability of LLM interventions to help users verify outputs and increase appropriate reliance.
2.2. LLMs as Decision Aids
Recent research has investigated the role of LLMs in human-LLM collaborations for applications such as writing (Liang et al., 2024), communication (Fu et al., 2024), coding (Kazemitabaar et al., 2023; Nam et al., 2024), education (Lo, 2023; Kumar et al., 2023b; Bo et al., 2024), fact-checking (Wolfe and Mitra, 2024; Kim et al., 2024a; Si et al., 2023), information search (Spatharioti et al., 2023; Salikutluk et al., 2023; Sharma et al., 2024), and other specialized domains (Choudhury and Shamszare, 2023; Cheong et al., 2024; Yang et al., 2023). In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations (Zhang, 2023). They can also offer antagonistic advice to induce the user into considering alternative options (Chiang et al., 2024; Ma et al., 2024a; Cai et al., 2024). Even when the user is just reading without explicitly searching for advice, the opinions embedded in the LLM outputs can be influence the user’s own opinions (Ferguson et al., 2023; Sharma et al., 2024; Krügel et al., 2023). In technical AI research, the quality of LLM responses has been improved through methods such as chain-of-thought with intermediate reasoning steps (Wei et al., 2022), debating amongst multiple LLMs (Khan et al., 2024; Kim et al., 2024a), and enforcing self-consistency in outputs (Wang et al., 2022; Becker and Soatto, 2024).
In many cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022). However, the question of how to induce appropriate reliance is still open. While using an LLM for information search can be just as efficient as a search engine, the user’s ability to distinguish incorrect information is diminished, resulting in a higher rate of over-reliance on incorrect information (Si et al., 2023; Xu et al., 2023). On the other hand, aversion to LLM advice has also been observed, resulting in users losing performance gains that they could have achieved through relying on the LLM (Zhang, 2023).
In addition to technical limitations of LLMs, human decision-makers are also imperfect and subject to cognitive biases (Bertrand et al., 2022). People may have unrealistic expectations of LLMs and use them inappropriately (Vafa et al., 2024), such as by assuming LLMs are skilled at math simply because they are computational models (Zhou et al., 2024). People also demonstrate divergent decision-making characteristics, with some preferring to meticulously optimize their decision while others take the satisficing or minimizing paths that allow them to bypass extensive thinking (Kaur et al., 2024; Misuraca et al., 2015). Those who minimize time spent on tasks may be more likely to over-rely on LLM advice, given its perceived completeness and relevance (Eigner and Händler, 2024).
2.3. Uncertainty and Reliance Calibration Techniques
Much attention has been placed on promoting transparency in LLMs, with the goal of communicating their capabilities or confidence to users (Liao and Vaughan, 2023; Yang et al., 2024; Lin et al., 2022). For example, Uncertainty Highlighting draws visual attention to uncertain tokens in the LLM’s outputs that the user may want to scrutinize (Do et al., 2024; Vasconcelos et al., 2023a), which have been found to reduce over-reliance in information search (Spatharioti et al., 2023), improve trust in answering questions (Do et al., 2024), and increase interpretability in AI-aided coding (Vasconcelos et al., 2023a). Other interventions that have been tested in human-LLM collaborations include contrastive explanations (Si et al., 2023), adding semantic expressions of uncertainty (Kim et al., 2024b; Lin et al., 2022; Steyvers et al., 2024), and confidence scores (Virk et al., 2024).
Complex LLM calibrations may carry the risk of being too cognitively overwhelming to understand—for example, to truly understand uncertainty highlighting, users must know that the highlighting reflects token-wise uncertainty rather than semantic uncertainty. Many people make satisficing decisions, where they only look at information to make ‘good enough’ decisions but don’t engage further, thus rendering explanations and interventions ineffective (Kaur et al., 2024; Eigner and Händler, 2024). Reliance Disclaimer is a low-effort intervention that can be statically applied to LLM outputs, such as through a persistent warning on the chat interface that encourages the user to verify the LLM’s answers. While extremely simple, it has been found to reduce students’ over-reliance on incorrect LLM guidance in a classroom setting (Bo et al., 2024).
Additionally, explainable AI (XAI) research has found it effective to focus on increasing the user’s cognitive engagement in the task, such as through simple cognitive forcing functions that add frictions and delays to the interface or removing direct recommendations (Buçinca et al., 2021; Gajos and Mamykina, 2022). Similar friction-based UI designs have been applied to the LLM space to modulate user behaviour (Collins et al., 2024). We take inspiration from the domain of LLMs for education, where LLMs are prompted to provide guidance or hints instead of direct answers (Kumar et al., 2023a; Ma et al., 2024b). Friction is introduced in the reliance process, where users have to cognitively engage to consider the recommended answer.
In this study, we implement three highly relevant interventions that induce cognitive engagement in different ways: Reliance Disclaimer, Uncertainty Highlighting, and Implicit Answer.
3. Methodology
We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We selected GPT-4o (https://openai.com/index/hello-gpt-4o/) as it is the most state-of-art LLM at the time of experimentation. in two question-answering tasks. To answer our research questions, we will compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the baseline LLM advice in the Control condition. In this section, we outline our choices of tasks, procedure, measures, and conditions.
3.1. Question-Answering Tasks
We gave the participants two separate question-answering tasks, which we carefully selected to meet the following desiderata:
D1
Accessible to a general audience. The tasks should not require any specialized knowledge or skills, so that they are relevant and accessible to a typical person.
D2
Difficult. The tasks should be sufficiently challenging, requiring a meaningful level of cognitive effort to answer correctly.
D3
Resistant to cheating. The tasks should not be straightforward to solve by using external assistance, such as a search engine.
Driven by the desiderata above, one of our central objectives was to select tasks that a typical person could reason about, even if they do not arrive at the correct answer. We therefore intentionally excluded factual questions characterized by domain-specific knowledge that individuals either possess or lack, and cannot conclude through reasoning alone. This approach emphasizes critical thinking over recall of factual information, and ensures that the average person has the potential to get the correct answer, as well as the ability to engage with LLM advice and reliance interventions. To increase generalizability, we selected two different question formats: multiple-choice, a common standard in human-AI collaboration research; and numerical estimation, which allows for a more granular measurement of the LLM’s influence on participants through the weight-of-advice (WoA). Below, we outline the tasks selected for each format.
LSAT Logical Reasoning. Based on preliminary results from piloting multiple-choice questions from several domains, we found that logical reasoning questions best balanced difficulty with accessibility. We selected LSAT logical reasoning questions as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’, for the first task. Each LSAT question has five options, of which exactly one is correct. We generated both correct and incorrect LLM advice for each question for use during the experiment. The questions and some sample LLM advice can be found in Figures A.1-A.2 in Appendix A. Correct answers were determined using LSAT answer keys.
While LSAT questions are highly detailed, they are designed to test for aptitude in law school rather than knowledge of the legal domain. As such, they do not contain undefined jargon, nor do they require domain knowledge (satisfying D1). In addition to reasoning, LSAT Logical Reasoning questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2). Furthermore, we reworded a question so that entering it into a search engine would not result in the correct answer (satisfying D3). Finally, they lend our study further external validity by emulating situations where people use LLMs to assist them with text comprehension or reasoning, in addition to people studying for tests such as the LSAT. While the MMLU dataset is another popular multiple-choice question benchmark for LLMs (Hendrycks et al., 2020), we did not find it suitable since it is skewed towards fact-based, know-it-or-you-don’t questions.
Numerical Estimation. While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a numerical estimation task adds a valuable new dimension to this research space. Among the various forms of numerical estimation, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar. We posit that this is a gamified task that is familiar to most people and does not require advanced mathematics (satisfying D1). Despite the simplicity of the estimation process, it is still challenging to achieve accurate results (satisfying D2). Lastly, since we took the pictures and many of the beans are occluded, it is impossible to use external assistance to “cheat” on this task (satisfying D3).
For the experiment, we captured photos of glass jars with manually-counted quantities of beans from both side perspectives and bottom perspectives. One set featured a smaller jar containing 403 chickpeas, while the other contained a larger jar with 856 jellybeans. The pictures and sample LLM advice can be found in Figures A.3-A.4 in Appendix A. Unlike in multiple-choice tasks, where advice can be objectively classified as good or bad, the goodness of the advice in the numerical estimation setting depends on the participant’s initial estimate. For example, while advice recommending a number that is far lower than the ground truth is objectively wrong, it is still better than an even lower participant estimate. To account for this variability, we provided five distinct levels of LLM advice relative to the ground truth quantity: very low, low, good, high, and very high. We chose these levels of advice based on results from initial pilot experiments to balance the proportion of bad advice with good advice. Visualizations of the advice levels with the distribution of participant guesses are shown in Figure C.7 in Appendix C.
3.2. Procedure
In the experiment, each participant was randomly assigned to one of four conditions and then answered two questions from the LSAT task and two questions from the Numerical Estimation task, randomized so that the LSAT phase and Numerical Estimation phase were equally likely to appear first. The orders of questions within each category were also randomized for each participant. Participants received pre-generated advice from GPT-4o. This advice was randomly sampled for each question, meaning participants could receive any mix of good or bad advice. After completing the tasks, participants completed a short post-survey. See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice from all four conditions.
To mimic the stakes and level of engagement typically involved in real-world decision-making scenarios where people might consult AI chatbots, we incorporated two key design elements into the study:
Two-Stage Answering: We aim to emulate contexts where the user is already engaged in the task before turning to LLM advice, such as a student who tries to solve the problem on their own first before checking with ChatGPT. Given that our participants are crowdworkers without prior context or investment in the tasks, we believed it is necessary to induce engagement. To address this, we require participants to answer each question on their own before seeing the LLM-generated advice, and then once again after they are given the advice. Studies that use one-stage questioning, where the LLM’s advice is presented alongside the initial question, reflect a context where the participant has less incentive to be engaged, which can lead to artificially high over-reliance rates.
Incentivization: Real-world decision-making involves balancing risk and reward. To simulate these stakes and induce incentive-compatible behaviour, we introduced a bonus, amounting to a maximum of 20% of the total participation pay. Participants earn rewards for correct answers, with two opportunities for bonus pay: one before, and one after receiving LLM-generated advice. In the LSAT questions, participants received a bonus if they selected the correct answer, while for the numerical estimation task, participants earned a bonus if they guessed within 10% of the true count.
3.3. Measures
Our research questions focus on reliance behaviour (RQ1) and confidence calibration (RQ2). We detail the equations for these categories of measurements, in addition to secondary measurements of task performance and time of engagement.
Reliance. Following existing literature on appropriate reliance in human-AI interactions, we use the two-stage question-answering task format where a participant first answers a given question on their own before receiving LLM advice and answering again. If the participant’s initial answer conflicts with the LLM’s advice, appropriate reliance requires the human to correctly determine whether the LLM’s advice is better or worse. We focus on this explicit decision-making step as the construct of interest for measuring reliance, where the user is presented with the choice to either a) change their answer to rely on the LLM’s advice, or b) disregard it and retain their initial answer. Thus in our analysis on reliance, we only consider the data instances where the participant’s initial answer differs from the LLM’s advice. See Figure 3 for an illustration of the reliance decisions pathways.
What does it mean to rely on LLM advice? In multiple-choice questions, we follow previous work by defining reliance as changing one’s answer to the LLM’s suggestion. In estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as . We further define the reliance outcomes as falling into one of four categories:
Positive LLM Reliance (PL): the participant changes their original answer to follow the LLM’s advice, and subsequently improves their performance.
Positive Self-Reliance (PS): the participant does not follow the LLM’s advice, and doing so would have worsened or failed to improve their performance.
Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance.
Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance.
PL and PS can be grouped together as Correct Reliance Behaviours, while NL and NS are Wrong Reliance Behaviours. NL represents over-reliance on LLM advice (the participant took the advice when it worsens or fails to improve improve their performance) and NS represents under-reliance (the participant didn’t take the advice but should have). We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extend to other real-world tasks. In contrast, the LSAT task lacks relative scoring across options, treating all incorrect answers as equally wrong; so relying on incorrect advice counts as over-reliance. This disambiguation is made since the LLM advice would not improve performance, so it would be preferable for the user to engage in the task and self-rely, allowing them to learn from the experience (Gajos and Mamykina, 2022). From these behaviour classifications, we compute the following appropriate reliance metrics:
Relative LLM Reliance (RLR) = Rate of appropriate reliance on correct LLM advice.
Relative Self-Reliance (RSR) = Rate of appropriate self-reliance when the LLM advice is wrong.
Appropriate Reliance Ratio (ARR) = Ratio of appropriate reliance on correct advice to over-reliance on wrong advice, where:
ARR ¡ 1: Higher rate of over-reliance than positive reliance.
ARR = 1: Random chance; LLM reliance rates are the same regardless of LLM correctness.
ARR ¿ 1: Higher rate of positive reliance than over-reliance.
We compute the metrics for each condition across participants, then calculate the effect size as the difference in outcomes between the control and each intervention. To assess statistical significance, we use permutation tests. For each planned comparison, we generate a distribution of effect sizes under the null hypothesis by randomly shuffling and reassigning data between the control and intervention groups. By comparing the observed effect size to this shuffled distribution, we estimate how likely it is to occur by chance, using a significance threshold of .
Confidence and Subjective Perceptions. For each question, we collect the participant’s confidence in the answer they gave prior to receiving LLM advice (Initial Confidence), and their confidence in the answer they gave after (Final Confidence) receiving the LLM-generated advice. Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence from the first to second stage. Well-calibrated confidence should result in lower confidence scores when participants make incorrect reliance decisions. We compare the change in confidence between correct and wrong reliance outcomes for each of the conditions, where statistical differences are tested with Mann-Whitney U-test for non-normal distributions with significance threshold .
At the end of the survey, participants provide additional ratings on a bipolar Likert scale of 1 (disagree) to 7 (agree) for the following subjective perception statements:
ChatGPT’s advice made me more confident.222This subjective confidence measure is separate from the question-level confidence measure (the actual target of RQ2).
ChatGPT’s advice seems accurate.
ChatGPT’s advice was helpful.
Task Performance. Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative task performance. However, we explore the initial error from participants independently answering questions, and compare this with their updated error after re-answering questions with LLM advice. We also compute a Best Possible Performance metric, which simulates the outcomes as if the human made the correct reliance decision in every single question, representing the best possible accuracy of a human-LLM team.
Time of Engagement. To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on answering each stage of question-answering. From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, and the time taken to answer the initial question (Relative Advice Time).
3.4. Experimental Conditions
In this section, we outline the implementation of our four experimental conditions, each representing different forms of LLM-generated advice. The interventions were chosen based on their relevance to the research community, their applicability across both multiple-choice and estimation tasks, and their diversity in design—incorporating both static and dynamic interventions, and both visual and semantic augmentations. Whenever possible, we used direct outputs generated through tailored prompts from the LLM with minimal manual adjustments. Figure 1 provides a visual overview of the advice conditions, with more details available in Figure 2(b).
Control. To generate the baseline advice, we input the questions into GPT-4o through OpenAI’s API and record the responses directly. Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to an explanation supporting the answer.
Reliance Disclaimer. The following static disclaimer statement is appended to all advice generated in Control: “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”. This text was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully. For consistency and to emulate real-world chatbot disclaimers like ChatGPT, we apply the same disclaimer to all advice. However, we acknowledge that the effectiveness of a static disclaimer may not persist in long-term deployment. Alarm fatigue describes the process of desensitization, where users learn to ignore unfounded warnings over time (Cvach, 2012). We hypothesize that disclaimers may be most effective for priming new users or in shorter interactions, like a tutoring chatbot that students only use occasionally. Some alternative ways to improve longevity of the effect are to dynamically insert disclaimers fluidly into the conversation—which makes them less repetitive and predictable (Bo et al., 2024), or to insert them only when error rates are particularly high (Cvach, 2012).
Uncertainty Highlighting. When generating the Control advice, we record the log probability of each token. Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a). To visually augment the baseline advice, we highlight the tokens with log probability under -1 in stronger red and indicate them to be uncertain, and the tokens under -0.5 in lighter pink and indicate them to be slightly uncertain. We selected these thresholds to balance highlighting uncertain text within the answer with minimizing visual overwhelm for the reader. We make manual adjustments to maintain coherence, such as extending highlighting to full words if they were only partially highlighted and removing highlighting on articles that do not impact semantic meaning. Participants were not told that the method of determining uncertainty is based on log probabilities. However, they were reminded of the definition of the highlighting colors in every survey page with the LLM advice.
Implicit Answer. This intervention semantically augments the content of the advice to imply the answer without writing it directly, requiring more engagement to extract the answer that the advice supports. For the Logical Reasoning questions, we append an instruction to the prompt to never explicitly state the answer’s text, and to use paraphrasing to generate an argument in support of the answer. For the Numerical Estimation questions, we manually removed the final calculation step from the advice generated in Control (as it was difficult to achieve this through prompting alone). In this way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
4. Results
We pre-registered our experimental design and analysis plan online333https://aspredicted.org/8BD_WWN, then ran the experiment. Here, we report the analyses described in Section 3 separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
4.1. Participants
We conducted a power analysis based on preliminary pilot data and found that approximately 200 samples per condition would suffice to achieve significance with an effect size of 0.4 for the ARR. Since we perform the analysis separately for LSAT and Numerical Estimation and each participant answers two questions, we determined that we needed 100 participants per condition.
We recruited 400 crowdworkers from Prolific, filtering for those who are based in the US, are fluent in English, have a task success rate of 98% or higher, and have completed at least 50 tasks. Following Prolific’s policy, we excluded participants who failed 2/3 attention checks, which was one participant from the Control () condition and two from the Uncertainty Highlighting condition (). For both of the other conditions, Reliance Disclaimer and Implicit Answer, all participants passed the attention checks. Across all conditions, the median completion time was 9 minutes and 44 seconds. Participants were compensated $1.28 USD as the base pay. They had the chance to earn up to $0.26 USD as a bonus for performance, with the mean earning being $0.07 USD. Based on Prolific’s demographics information about our participants, the participants are 56.3% female with an average age of 43.5 years and are all located in the US. Furthermore, they are well-distributed in their usage of AI chatbots, with 18.9% citing daily use, 33.5% citing weekly use, 17.9% citing monthly use, and the remaining 29.7% citing rare or no usage.
4.2. Task Performance
We preface our main discussion of reliance with an examination of task performance. We emphasize that since we are interested in how people react to both good and bad LLM advice, we sample these at equal rates. Our LLM performance therefore does not necessarily reflect the actual performance of GPT-4o on our tasks and should not be taken as a measure of realistic human-LLM collaborative decision-making.
LSAT. Due to the high difficulty of the LSAT questions, participants were initially correct in only 36.0% of question instances, not far from the rate of random guessing (20%). Participants then either received correct or wrong LLM advice, equivalent to a 50% accuracy rate. Their second-stage (LLM-assisted) accuracy rate was 46.1%. Figure 4 shows transition matrices of the first- and second-stage performance outcomes in each condition. In the performance outcomes, there are no significant differences between Control and any of the interventions as evaluated by a test ( for all three interventions). Based on the collaborative performance, it is evident that while the LLM improved overall performance, the human-AI team does not perform optimally. In each question, if the human-LLM team relied on the team member that got the correct answer, then the Best Possible Performance would reach 62.1% accuracy (not 100%, because sometimes both the human and the LLM answered incorrectly). The gap between optimal and actual performance represents the performance lost to inappropriate reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
Numerical Estimation. While estimation may be more approachable, it is still a difficult task. The mean absolute percentage error relative to the ground truth is 45.7% across all conditions. With advice, the participants do not significantly improve, answering with a similar absolute error rate of 47.0%—this is unsurprising, since poor LLM advice was given as frequently as good advice. However, in terms of the mean (not absolute) percentage error, the participants initially underestimate the counts, averaging -20.2% relative to the ground truth. With the advice, the gap closes to just -5.6% (below the ground truth). The full distributions, which include both positive and negative percent errors, are shown by condition in Figure 5(a). There are no significant performance differences between conditions (Mann-Whitney U test for all three interventions). The optimal human-LLM team outcome can be simulated by submitting the estimation of the team member who is closest to ground truth. In this hypothetical Best Possible Performance, the mean of the absolute error would be 29.0%. That the observed error was much higher indicates people both over- and under-relied on LLM advice.
While unrelated to task performance, we further compute how much the participants relied on the advice for additional insights on their behaviour. The weight-of-advice taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions. The distribution of WoA tends to be bimodal, with participants in Control, Reliance Disclaimer, and Uncertainty Highlighting predominantly favouring adoption of the LLM advice. Since Implicit Answer did not provide the final recommendation, and instead left it to the participants to apply the mathematical hints, the WoA distribution is more flattened, even with values in the negatives or exceeding 1. In our WoA analysis, we drop outlier values over a magnitude of 5 (such as switching from an initial estimate of 600 to a final estimate 4480 when the LLM advice is 544, which constitutes a WoA of -69.3). By the Mann-Whitney U test, Reliance Disclaimer’s WoA () and Implicit Answer’s WoA () are significantly smaller than Control’s WoA.
4.3. Reliance Behaviour
Following the methods of previous work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases. For Numerical Estimation, we do not use a strict definition of agreement, as it is unlikely for a participant’s initial estimate to exactly match the LLM advice. Instead, we adopt a heuristic of perceived agreement. For example, a participant may perceive a guess of 700 to be “pretty much the same” as an advice of 750. We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.444We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small. To address this, we include an additional 50 padding around the advice. As a robustness check, we also conduct reliance analysis using a 5% agreement boundary without additional padding and find similar significance trends to those reported. After excluding the instances of agreement, 87.1% of samples in the LSAT task and 88.8% in Numerical Estimation are retained.
To answer RQ1, we compute the Relative LLM Reliance (RLR), Relative Self-Reliance (RSR), and Appropriate Reliance Ratio (ARR) across all data instances collected per task in each condition. For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is bad. ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM (less over-reliance and less under-reliance, or at least a positive trade-off such as much less over-reliance and only a small increase in under-reliance). In terms of reliance outcomes, over-reliance maps to low RSR (i.e. relying on the LLM when the advice is bad), under-reliance maps to low RLR (i.e. not relying on the LLM when the advice is good), and appropriate reliance maps to when both are high. Figure 6(a), reconstructed based on Figure 1 from (Schemmer et al., 2022), provides a visual reference for the relationship between reliance outcomes and RLR and RSR values. Table 1 displays all the metrics calculated for each condition, separately for the LSAT and Numerical Estimation tasks. For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control.
| Control | Reliance Disclaimer | Uncertainty Highlighting | Implicit Answer | ||
| LSAT | RLR | 0.66 | 0.71, p=.2 | 0.52, p=.1 | 0.58, p=.2 |
| RSR | 0.44 | 0.56, p=.05 | 0.62, p=.01 (+) | 0.63, p=.002 (+) | |
| ARR | 1.17 | 1.62, p=.04 (+) | 1.35, p=.2 | 1.54, p=.09 | |
| Numerical Estimation (WoA >0.01) | RLR | 0.91 | 0.91, p=.4 | 0.93, p=.3 | 0.67, p<.001 (–) |
| RSR | 0.15 | 0.22, p=.1 | 0.23, p=.1 | 0.43, p=.004 (+) | |
| ARR | 1.07 | 1.17, p=.2 | 1.20, p=.1 | 1.18, r=.2 | |
| Numerical Estimation (WoA >0.5) | RLR | 0.8 | 0.74, p=.2 | 0.8, p=.5 | 0.45, p=.17 |
| RSR | 0.39 | 0.47, p=.1 | 0.42, p=.3 | 0.55, p=.007 (+) | |
| ARR | 1.32 | 1.38, p=.4 | 1.37, p=.4 | 1.01, p=.1 | |
| Numerical Estimation (WoA >0.99) | RLR | 0.47 | 0.39, p=.2 | 0.3, p=.02 (–) | 0.34, p=.05 |
| RSR | 0.68 | 0.79, p=.02 (+) | 0.77, p=.05 | 0.67, p=.5 | |
| ARR | 1.45 | 1.91, p=.2 | 1.34, p=.4 | 1.03, p=.1 |
LSAT. The visual representations of the metrics and their significances for LSAT are plotted in Figure 6(b), where Control is plotted in gray and we compare the interventions against it. We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance (i.e. it is higher and further to the right than Control). Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance, but at the cost of shifting reliance away from the LLM when following it would be appropriate (although this decrease is not statistically significant)—these interventions tend to reduce LLM reliance overall, both when the advice is bad and when it is good. Both interventions have higher ARR values than Control, but the difference is not significant at our power level. These results indicate that a static, persistent disclaimer message can induce appropriate reliance more effectively than more specialized methods.
Numerical Estimation. This task requires more care to analyze, as WoA is a continuous rather than binary measure of reliance. As such, we compute reliance behaviours at several different WoA thresholds:
WoA ¿ 0.01: Any adjustment towards the LLM’s advice is counted as reliance (highest sensitivity). Under this definition, 79.7% of trials count as reliant behaviour.
WoA ¿ 0.5: Adjustments must be made over 50% towards the LLM’s advice to count as reliance (moderate sensitivity). Under this definition, 60.7% of trials count as reliant behaviour.
WoA ¿ 0.99: Adjustment towards the LLM’s advice must be nearly total to count as reliance (lowest sensitivity). Under this definition, 32.7% of trials count as reliant behaviour.
Values and significance for RLR, RSS, and ARR are expressed at these three WoA threshold levels in both Table 1 and Figure 6(c). Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance. Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 and 0.5). At the lowest WoA threshold, this also increases under-reliance—which is undesirable. This could be due to being forced to perform simple math calculations to see the advised number—which people may not do if they are satisfied with their own answer, or believe that extracting the advice requires more effort than the reward is worth. Uncertainty Highlighting also causes significant decrease in RLR at high WoA thresholds, indicating that people are hesitant to fully use the advice. This corroborates Figure 5(b), which shows that Uncertainty Highlighting and Reliance Disclaimer both have reduced density of WoA values close to 1 compared to Control.
Another key difference between the Numerical Estimation and LSAT tasks is the length and detail of the advice. A higher barrier to cognitive engagement may lead to hetereogenous outcomes in how participants approach the task. For instance, some participants might adopt the advice as a shortcut to make easy, satisficing decisions, while others might disregard the advice due to its excessive length and complexity (Eigner and Händler, 2024). The discrepancy in results between our two tasks also warns that results of interventions may not generalize beyond the specific tasks they are tested on, emphasizing the need for thorough validation across diverse tasks before broader deployment.
4.4. Confidence Calibration and Perceptions
To address RQ2, we examine participant confidence in their pre-advice answer and their post-advice answer for each question, along with several post-experiment Likert scale ratings of their subjective perceptions. We compare the changes in confidence between cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (over-reliance). Unless otherwise indicated, significance is calculated within each condition using the Mann-Whitney U test with as the significance threshold. Error bars in the bar plots of Figures 7-10 represent standard errors.
LSAT. Figure 7(a) shows participants’ LSAT question confidence changes for correct and wrong reliance behaviours. The most salient takeaway is that on-aggregate across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions (). Within each condition, Uncertainty Highlighting is the only one where this observed effect is significant (), which indicates poor confidence calibration since it is undesirable for people to be more confident in wrong decisions.
As an extension, Figure 7(b) specifically focuses on cases where people chose to rely on the LLM. Here, the confidence scores tend to increase more as people trusted the LLMs enough to use them. We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is significantly lower when the participant over-relied than when they relied appropriately (). Though not significant, it is interesting to observe that the opposite trend can be demonstrated in the other interventions, where confidence increases more in cases of over-reliance. This indicates that reliance interventions that operate by moderating users’ confidence in the LLM may inadvertently miscalibrate confidence with respect to appropriate reliance.
Numerical Estimation. For this task, we repeat the confidence analyses and show them in Figure 8. For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.8-D.9). Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes. However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs, which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
Overall Post-Perception. Lastly, Figure 9 shows the mean Likert ratings per perceptions category by condition. The full subjective perception statements are shown in the figure’s labels (1=disagree, 7=agree). We calculate significance between the Control and the interventions using the Mann-Whitney U test with as threshold. Although all LLM conditions scored positively across the three categories, Control scored highest in subjective perceptions compared to all other the interventions. In particular, Uncertainty Highlighting is the only intervention that had significantly lower ratings in all categories compared to Control, indicating that our implementation may have induced too much uncertainty in the users. An alternative approach to moderate this effect may be to highlight highly certain tokens in green in addition to uncertain tokens in red. In addition, we find that the frequency of LLM usage correlates positively with all three subjective perception measures, but not with reliance behaviours – see Appendix E for details. The main implication here is that uncertainty calibration methods can have the undesirable outcome of inducing negative perceptions of accuracy and usefulness of the LLM, which increases overall algorithm aversion against LLM advice, rather than appropriate reliance.
4.5. Time of Engagement
We compare the relative time of engagement with given LLM advice by taking the ratio of time spent in the second stage to the first stage. Time is correlated to cognitive engagement, and previous studies have found that introducing slight temporal frictions like waiting for delayed AI advice has potential to affect user behaviour (Buçinca et al., 2021; Collins et al., 2024). Figure 10 shows the time ratio by condition for both the LSAT (Figure 10(a)) and Numerical Estimation (Figure 10(b)) tasks, where significance is indicated by Mann-Whitney U test with . Overall, participants spent more time in the Numerical Estimation question, with the average relative time ratio exceeding 1, potentially due to the average length of the advice being longer. On the other hand, the time ratio is well below 1 in LSAT.
In both tasks, the order of relative time spent on the advice roughly matches the effort required to extract the answer from the advice. Reliance Disclaimer is closest to matching the time ratio of Control while Implicit Answer takes much longer, reaching nearly double the relative time in Numerical Estimation. Uncertainty Highlighting is somewhere in between, but also with significant increase in time over Control. Taken with the results on the efficacy of Reliance Disclaimer, our results indicate that longer engagement time does not necessarily correlate with reliance outcomes. The easier, more digestible intervention had the most tangible effect on appropriate reliance. Perhaps with more training and rigour, especially for Uncertainty Highlighting, participants can become more proficient at interpreting advice with more complex uncertainty calibration.
5. Discussion
We now present the key findings to our research questions on appropriate reliance and confidence calibration. We then extend our findings to broader implications for design and for evaluating reliance on LLMs, and discuss the limitations of our experiments.
5.1. Key Findings
Cognitive Forcing Designs Reduce Over-Reliance, but May Not Improve Appropriate Reliance. In our study, we investigated the effectiveness of three interventions on LLM reliance in two diverse and challenging tasks. We use measures of inappropriate reliance, separating over-relying on bad advice from under-relying on helpful advice, to uncover mechanisms of sub-optimal human-LLM team performance. Our findings suggest that cognitive forcing designs in the form of confidence calibration (highlighting uncertain tokens) and usability friction (adding disclaimers and removing direct answers) can help people shift away instinctively relying on the LLM, but improving appropriate advice usage is a harder challenge.
While the effect of each intervention varied by the task, all three achieved a reduction in over-reliance in at least some experiment configurations. However, the only intervention to reduce both over-reliance and under-reliance is Reliance Disclaimer in the LSAT logical reasoning task. The Numerical Estimation task captured reliance with more granularity, which provided insight on how the interventions induced under-reliance differently. Our results suggest that the added friction of calculating the answer in Implicit Answer caused people to disproportionately discard the LLM advice (not rely at all), while the confidence calibration in Uncertainty Highlighting caused people to hesitate to fully follow advice (rely, but not fully).
The increase in advice under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b). This suggests that for the LSAT task, static Reliance Disclaimers may have added just enough friction to dissuade habitual over-reliance. However, advice with other interventions, like Implicit Answer in Numerical Estimation, may have been perceived as requiring more effort to use than it was worth. Future work should aim to quantify the comprehension difficulty of the interventions and the tasks to understand how they relate to reliance behaviours. In implementing interventions, designers should consider the targeted use case and develop cognitive forcing techniques that do not increase burden substantially.
Confidence When Using LLMs is Poorly Calibrated. We examine how participants’ confidence in their answers change after seeing the LLM’s advice. Ideally, even with incorrect reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice. However, we largely observe the opposite or inconclusive effects, where wrong reliance results in higher gains in confidence. This raises particularly problematic implications, suggesting that individuals were overall uncalibrated when receiving assistance from LLMs, corroborating prior findings (Zhang, 2023; Si et al., 2023)
There were a few areas where the interventions had success. In the Numerical Estimation task, Implicit Answer reduced confidence during under-reliance, while in the LSAT task, Reliance Disclaimer lowered confidence gains during over-reliance, which can be exploited as feedback mechanisms for users to learn from. Both methods present different pathways using which confidence can be influenced by modifying LLM output, but further work is needed to achieve generalizable findings. As shown in prior works in human-AI collaboration, self-confidence and advice agreement rates have nuanced relationships with reliance on advice (Lu and Yin, 2021), which can be further investigated in the context of human-LLM collaboration.
We also find perceptions of helpfulness were negatively impacted by all three interventions, while perceived self-confidence and advice accuracy were only reduced by Uncertainty Highlighting. Confidence calibration interventions may risk increasing a user’s aversion towards LLMs rather than encouraging appropriate evaluation if overall performance is perceived as unacceptably low, or if the intervention makes the advice difficult to understand and use. Further work can explore qualitative factors of what participants disliked or found more cumbersome, and how their perception of their internal decision-making process was affected.
5.2. Broader Implications
With the integration of Copilot in consumer PCs, the widespread use of LLMs in customer service and information retrieval, and advances in Generative AI for creative work, interactions between end users and these tools are becoming nearly unavoidable. As Generative AI advances rapidly, consumers’ literacy in its appropriate use is likely to lag behind, raising concerns about how to ensure proper use of the many AI-powered tools and products being introduced (Annapureddy et al., 2024). In our study, we show that wrongful reliance is frequent and users are not calibrated to be aware of their behaviour. Interventions that reduce over-reliance can be have undesired effects of increasing under-reliance, and require thorough validation in human user studies. We highlight the need to extend the appropriate reliance metrics to realistic usage scenarios.
Generalization of the Appropriate Reliance Metrics. Through our experiments, we highlight the flexibility of the Relative LLM Reliance, Relative Self-Reliance, and Appropriate Reliance Ratio metrics in benchmarking interventions across two diverse task formats. While prior works have focused on binary or multiple choice answers, we demonstrate the suitability of the metrics for numerical targets, which opens up their applicability to additional real-world tasks. These measures enable more nuanced analyses of suboptimal human-LLM collaboration beyond task performance alone, and we recommend their incorporation into future empirical research on LLM usage. The metrics can be generalized to any LLM-assisted task where there is a deterministic notion of optimality (i.e. some outcomes are preferred over others), such as code compilation success for programming, accuracy of arguments in writing persuasive essays, and requirement satisfaction in travel planning. Extending experiments to more open-ended tasks requires further nuance to faithfully capture agreement, reliance, and correctness. However, we theorize some degree of generalization in our high level findings—that cognitive forcing interventions will cause people to approach reliance on LLM content more cautiously.
LLM Design Trade-Offs Based on the Context of Use. The question, “How can interventions be designed successfully?”, still remains open. We uncover complex heterogeneity in how interventions performed across different tasks, indicating that optimal designs are likely context-dependent rather than universal. Our work motivates further investigation into the theoretical foundations linking individual factors, tasks characteristics, and LLMs’ response styles, and towards developing a reliable empirical framework to evaluate novel interventions rigorously. Additionally, as our results suggest a trade-off between reducing over-reliance and increasing under-reliance, it is therefore vital that practitioners consider the context of use of the LLM tool. This involves a closer evaluation of the stakeholders, users, and risks involved in the decision-making tasks (Liao et al., 2022).
In medical or legal decision-making, the heightened risk requires accountability, and errors from over-reliance may cause extreme damage if unchecked (Cheong et al., 2024; Yang et al., 2023). In creative tasks where the diversity of humans is valued, over-reliance may introduce an undesirable degree of homogeneity (Kumar et al., 2024). Within the education space, over-reliance brings about another issue: the unintended reduction of skill development due to relying on the LLM as a cognitive crutch (Zhai et al., 2024). As AI capabilities advance and eventually eclipse human capabilities, avoiding under-reliance may be preferred in lieu of aversion. However, scaling ability requires scaling oversight, representing further challenges in human-AI collaboration that require carefully calibrated trust and reliance (Bowman et al., 2022).
Human Validation for Technical Development. Ensuring the success of human-AI interactions requires bi-directional design and validation from both AI and HCI researchers (Shen et al., 2024; Vodrahalli et al., 2022). Technical AI research has progressed significantly towards understanding LLMs’ reasoning pathways, reducing hallucinations, and calibrating the confidence of outputs (Ayyamperumal and Ge, 2024; Feng et al., 2024; Lin et al., 2022; Zhang et al., 2023). However, such works often rely on automating evaluations by using existing benchmarks or artificially simulating users, considering only theoretical design desiderata and bypassing human evaluation. These evaluation methods do not ascertain whether these methods will have the desired effects when deployed to real users. For example, we showed that highlighting log probabilities to calibrate confidence did not have the intended effect on humans, who may have found the added information too confusing to optimally leverage. Not only did reliance and performance outcomes lack improvement, we also found that participants’ confidence changes were poorly calibrated, with a significant increase in post-task confidence when LLM advice is used inappropriately. We therefore encourage the broader research community, particularly the technical AI community, to capture multi-dimensional measures from real users.
5.3. Limitations
Overall, our results would be further strengthened by future research that explores LLM reliance in more naturalistic and long-term settings, offering a clearer picture of how users assess and use model outputs in the real world. Incorporating qualitative measures can also provided more nuanced feedback on how the participants regarded the advice and made choices.
Engagement of Participants in Tasks. The generalizability of our results is primarily limited by our design of a lab experiment with crowdworkers recruited from Prolific. Since participants knew that they were participating in an experiment, they may have inferred that they were being evaluated on LLM reliance and become more cautious about their reliance behaviours. Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to “hedge”; to maximize their chances of earning some bonus payment by choosing different answers at each stage. Our two-stage question answering set-up also emulates a specific type of decision-making where users are already engaged in the task. In fact, having users answer questions independently before advice can be considered a form of cognitive forcing even without additional interventions (Ashktorab et al., 2024). As LLM usage becomes more common, it is expected that many users will bypass this critical engagement and rely on LLMs by default, which can reflect in reliance outcomes differently.
External Validity of the Tasks. Our chosen tasks do not perfectly translate to the usage contexts for which people consult LLMs in the real world. While multiple-choice questions provided clean correctness and reliance metrics, real-world tasks predominantly consist of open-ended questions and generative tasks (such as writing and coding). Knowledge-based questions were avoided in this study to ensure that the questions would be accessible to participants; however, real questions posed to LLMs frequently require synthesis of more complex ideas and knowledge bases, which may be more susceptible to factual hallucinations.
Realism of Interacting with the LLM. Due to the nature of the survey study, our participants were unable to interactively ask follow-up questions or use sanity checks to verify the model’s outputs and prevent consequential errors. To ensure clean reliance metrics, they saw fixed advice and were asked not to consult external resources. Although we used direct outputs generated through tailored LLM prompts whenever possible, we did apply manual adjustments in some interventions and artificially manipulated the recommendation in the LLM’s advice. The distribution of advice quality was held constant regardless of the LLM’s true tendencies when responding to the question to prevent question difficulty from being a confound. We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude555https://claude.ai/, Copilot666https://copilot.microsoft.com/, Gemini777https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour. We also did not explore varying the system prompts to steer the LLM into embodying different personas. Furthermore, the framing of the LLM (e.g., ChatGPT vs. a generic chatbot) can prime user expectations of the quality and accuracy of the responses.
Selection of Interventions. To operationalize our research questions, we made certain design choices in the implementation of the interventions. For example, we used a naive method of highlighting based on token generation probability, while prior works have identified that likelihood of editing is a better indicator (Vasconcelos et al., 2023a). Our evaluation of the interventions also did not include any extensive training or longer-term follow-ups. As such, the impacts of each intervention in our studies are not necessarily representative of how they would impact reliance if optimally implemented in the real world. In particular, it is not clear if static disclaimers would persistently encourage appropriate reliance in deployment where users may experience alarm fatigue (Cvach, 2012), whereas dynamic disclaimers may have a longer impact (Bo et al., 2024). We explored a very small space of potential interventions, and encourage further human-centered evaluation of other uncertainty calibration techniques (e.g., confidence score (Virk et al., 2024), semantic expressions of uncertainty (Kim et al., 2024b; Lin et al., 2022; Steyvers et al., 2024), and contrastive explanations (Si et al., 2023)) and setting appropriate expectations of the LLM (e.g., model report card (Yang et al., 2024) and usage tutorial (Chiang and Yin, 2022)).
6. Conclusion
We measured the effects of three different LLM interventions on users’ reliance behaviours and confidence levels in two difficult LLM-assisted question-answering tasks. We find that adding a persistent disclaimer is the most effective at improving appropriate reliance (in the LSAT task) without significantly increasing the time needed to engage with the LLM advice, while both highlighting uncertain tokens and removing direct answers only reduced over-reliance and increased the engagement time. In addition, people’s confidence levels were poorly calibrated, generally exhibiting higher confidence increases when they made the wrong reliance decision. Based on our results, we call for the following: rigorous and human-centered evaluation of novel LLM reliance interventions; consideration for trade-offs between over- and under-reliance based on the context of use; and future research in naturalistic human-LLM collaboration tasks.
Acknowledgements.
We thanks Nicholas Carlini for helpful discussions during the early phase of the project. We further thank Lev McKinney, Lillio Mok, Harsh Kumar, Nakwon Rim, Paula Aoyagui, Zixin Zhao, and members of the UofT Computational Social Science Lab for feedback on the experimental methods and the draft of the manuscript.References
- (1)
- Annapureddy et al. (2024) Ravinithesh Annapureddy, Alessandro Fornaroli, and Daniel Gatica-Perez. 2024. Generative AI literacy: Twelve defining competencies. Digital Government: Research and Practice (2024).
- Ashktorab et al. (2024) Zahra Ashktorab, Qian Pan, Werner Geyer, Michael Desmond, Marina Danilevsky, James M Johnson, Casey Dugan, and Michelle Bachman. 2024. Emerging Reliance Behaviors in Human-AI Text Generation: Hallucinations, Data Quality Assessment, and Cognitive Forcing Functions. arXiv preprint arXiv:2409.08937 (2024).
- Ayyamperumal and Ge (2024) Suriya Ganesh Ayyamperumal and Limin Ge. 2024. Current state of LLM Risks and AI Guardrails. arXiv preprint arXiv:2406.12934 (2024).
- Bansal et al. (2021) Gagan Bansal, Tongshuang Wu, Joyce Zhou, Raymond Fok, Besmira Nushi, Ece Kamar, Marco Tulio Ribeiro, and Daniel Weld. 2021. Does the whole exceed its parts? The effect of AI explanations on complementary team performance. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–16.
- Becker and Soatto (2024) Evan Becker and Stefano Soatto. 2024. Cycles of Thought: Measuring LLM Confidence through Stable Explanations. arXiv preprint arXiv:2406.03441 (2024).
- Bertrand et al. (2022) Astrid Bertrand, Rafik Belloum, James R Eagan, and Winston Maxwell. 2022. How cognitive biases affect XAI-assisted decision-making: A systematic review. In Proceedings of the 2022 AAAI/ACM Conference on AI, Ethics, and Society. 78–91.
- Bo et al. (2024) Jessica Bo, Harsh Kumar, Michael Liut, and Ashton Anderson. 2024. Disclosures & Disclaimers: Investigating the Impact of Transparency Disclosures and Reliability Disclaimers on Learner-LLM Interactions. Available at SSRN (2024).
- Bowman et al. (2022) Samuel R Bowman, Jeeyoon Hyun, Ethan Perez, Edwin Chen, Craig Pettit, Scott Heiner, Kamilė Lukošiūtė, Amanda Askell, Andy Jones, Anna Chen, et al. 2022. Measuring progress on scalable oversight for large language models. arXiv preprint arXiv:2211.03540 (2022).
- Buçinca et al. (2021) Zana Buçinca, Maja Barbara Malaya, and Krzysztof Z Gajos. 2021. To trust or to think: cognitive forcing functions can reduce overreliance on AI in AI-assisted decision-making. Proceedings of the ACM on Human-computer Interaction 5, CSCW1 (2021), 1–21.
- Burton et al. (2020) Jason W Burton, Mari-Klara Stein, and Tina Blegind Jensen. 2020. A systematic review of algorithm aversion in augmented decision making. Journal of behavioral decision making 33, 2 (2020), 220–239.
- Cai et al. (2024) Alice Cai, Ian Arawjo, and Elena L Glassman. 2024. Antagonistic AI. arXiv preprint arXiv:2402.07350 (2024).
- Cao and Huang (2022) Shiye Cao and Chien-Ming Huang. 2022. Understanding user reliance on AI in assisted decision-making. Proceedings of the ACM on Human-Computer Interaction 6, CSCW2 (2022), 1–23.
- Chen and Shu (2023) Canyu Chen and Kai Shu. 2023. Combating misinformation in the age of llms: Opportunities and challenges. AI Magazine (2023).
- Chen et al. (2023) Valerie Chen, Q Vera Liao, Jennifer Wortman Vaughan, and Gagan Bansal. 2023. Understanding the role of human intuition on reliance in human-AI decision-making with explanations. Proceedings of the ACM on Human-computer Interaction 7, CSCW2 (2023), 1–32.
- Cheong et al. (2024) Inyoung Cheong, King Xia, KJ Kevin Feng, Quan Ze Chen, and Amy X Zhang. 2024. (A) I Am Not a Lawyer, But…: Engaging Legal Experts towards Responsible LLM Policies for Legal Advice. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. 2454–2469.
- Chiang et al. (2024) Chun-Wei Chiang, Zhuoran Lu, Zhuoyan Li, and Ming Yin. 2024. Enhancing AI-Assisted Group Decision Making through LLM-Powered Devil’s Advocate. In Proceedings of the 29th International Conference on Intelligent User Interfaces. 103–119.
- Chiang and Yin (2022) Chun-Wei Chiang and Ming Yin. 2022. Exploring the effects of machine learning literacy interventions on laypeople’s reliance on machine learning models. In Proceedings of the 27th International Conference on Intelligent User Interfaces. 148–161.
- Choudhury and Shamszare (2023) Avishek Choudhury and Hamid Shamszare. 2023. Investigating the impact of user trust on the adoption and use of ChatGPT: survey analysis. Journal of Medical Internet Research 25 (2023), e47184.
- Collins et al. (2024) Katherine M Collins, Valerie Chen, Ilia Sucholutsky, Hannah Rose Kirk, Malak Sadek, Holli Sargeant, Ameet Talwalkar, Adrian Weller, and Umang Bhatt. 2024. Modulating Language Model Experiences through Frictions. arXiv preprint arXiv:2407.12804 (2024).
- Cvach (2012) Maria Cvach. 2012. Monitor alarm fatigue: an integrative review. Biomedical instrumentation & technology 46, 4 (2012), 268–277.
- Dietvorst et al. (2015) Berkeley J Dietvorst, Joseph P Simmons, and Cade Massey. 2015. Algorithm aversion: people erroneously avoid algorithms after seeing them err. Journal of experimental psychology: General 144, 1 (2015), 114.
- Do et al. (2024) Hyo Jin Do, Rachel Ostrand, Justin D Weisz, Casey Dugan, Prasanna Sattigeri, Dennis Wei, Keerthiram Murugesan, and Werner Geyer. 2024. Facilitating Human-LLM Collaboration through Factuality Scores and Source Attributions. arXiv preprint arXiv:2405.20434 (2024).
- Eigner and Händler (2024) Eva Eigner and Thorsten Händler. 2024. Determinants of llm-assisted decision-making. arXiv preprint arXiv:2402.17385 (2024).
- Feng et al. (2024) Shangbin Feng, Weijia Shi, Yike Wang, Wenxuan Ding, Vidhisha Balachandran, and Yulia Tsvetkov. 2024. Don’t Hallucinate, Abstain: Identifying LLM Knowledge Gaps via Multi-LLM Collaboration. arXiv preprint arXiv:2402.00367 (2024).
- Ferguson et al. (2023) Sharon A Ferguson, Paula Akemi Aoyagui, and Anastasia Kuzminykh. 2023. Something Borrowed: Exploring the Influence of AI-Generated Explanation Text on the Composition of Human Explanations. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. 1–7.
- Fok and Weld (2023) Raymond Fok and Daniel S Weld. 2023. In search of verifiability: Explanations rarely enable complementary performance in AI-advised decision making. AI Magazine (2023).
- Fu et al. (2024) Yue Fu, Sami Foell, Xuhai Xu, and Alexis Hiniker. 2024. From Text to Self: Users’ Perception of AIMC Tools on Interpersonal Communication and Self. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–17.
- Gajos and Mamykina (2022) Krzysztof Z Gajos and Lena Mamykina. 2022. Do people engage cognitively with AI? Impact of AI assistance on incidental learning. In Proceedings of the 27th International Conference on Intelligent User Interfaces. 794–806.
- Guo et al. (2024) Ziyang Guo, Yifan Wu, Jason D Hartline, and Jessica Hullman. 2024. A Decision Theoretic Framework for Measuring AI Reliance. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. 221–236.
- He et al. (2023) Gaole He, Lucie Kuiper, and Ujwal Gadiraju. 2023. Knowing about knowing: An illusion of human competence can hinder appropriate reliance on AI systems. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–18.
- Hendrycks et al. (2020) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2020. Measuring massive multitask language understanding. arXiv preprint arXiv:2009.03300 (2020).
- Inkpen et al. (2023) Kori Inkpen, Shreya Chappidi, Keri Mallari, Besmira Nushi, Divya Ramesh, Pietro Michelucci, Vani Mandava, Libuše Hannah Vepřek, and Gabrielle Quinn. 2023. Advancing human-AI complementarity: The impact of user expertise and algorithmic tuning on joint decision making. ACM Transactions on Computer-Human Interaction 30, 5 (2023), 1–29.
- Kabir et al. (2024) Samia Kabir, David N Udo-Imeh, Bonan Kou, and Tianyi Zhang. 2024. Is stack overflow obsolete? an empirical study of the characteristics of chatgpt answers to stack overflow questions. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–17.
- Kaur et al. (2024) Harmanpreet Kaur, Matthew R Conrad, Davis Rule, Cliff Lampe, and Eric Gilbert. 2024. Interpretability Gone Bad: The Role of Bounded Rationality in How Practitioners Understand Machine Learning. Proceedings of the ACM on Human-Computer Interaction 8, CSCW1 (2024), 1–34.
- Kazemitabaar et al. (2023) Majeed Kazemitabaar, Justin Chow, Carl Ka To Ma, Barbara J Ericson, David Weintrop, and Tovi Grossman. 2023. Studying the effect of AI code generators on supporting novice learners in introductory programming. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–23.
- Khan et al. (2024) Akbir Khan, John Hughes, Dan Valentine, Laura Ruis, Kshitij Sachan, Ansh Radhakrishnan, Edward Grefenstette, Samuel R Bowman, Tim Rocktäschel, and Ethan Perez. 2024. Debating with more persuasive llms leads to more truthful answers. arXiv preprint arXiv:2402.06782 (2024).
- Kim et al. (2023) Antino Kim, Mochen Yang, and Jingjing Zhang. 2023. When algorithms err: Differential impact of early vs. late errors on users’ reliance on algorithms. ACM Transactions on Computer-Human Interaction 30, 1 (2023), 1–36.
- Kim et al. (2024a) Kyungha Kim, Sangyun Lee, Kung-Hsiang Huang, Hou Pong Chan, Manling Li, and Heng Ji. 2024a. Can LLMs Produce Faithful Explanations For Fact-checking? Towards Faithful Explainable Fact-Checking via Multi-Agent Debate. arXiv preprint arXiv:2402.07401 (2024).
- Kim et al. (2024b) Sunnie SY Kim, Q Vera Liao, Mihaela Vorvoreanu, Stephanie Ballard, and Jennifer Wortman Vaughan. 2024b. ” I’m Not Sure, But…”: Examining the Impact of Large Language Models’ Uncertainty Expression on User Reliance and Trust. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. 822–835.
- Kleinberg et al. (2018) Jon Kleinberg, Himabindu Lakkaraju, Jure Leskovec, Jens Ludwig, and Sendhil Mullainathan. 2018. Human decisions and machine predictions. The quarterly journal of economics 133, 1 (2018), 237–293.
- Krügel et al. (2023) Sebastian Krügel, Andreas Ostermaier, and Matthias Uhl. 2023. ChatGPT’s inconsistent moral advice influences users’ judgment. Scientific Reports 13, 1 (2023), 4569.
- Kumar et al. (2023a) Harsh Kumar, Ilya Musabirov, Mohi Reza, Jiakai Shi, Anastasia Kuzminykh, Joseph Jay Williams, and Michael Liut. 2023a. Impact of guidance and interaction strategies for LLM use on Learner Performance and perception. arXiv preprint arXiv:2310.13712 (2023).
- Kumar et al. (2023b) Harsh Kumar, David M Rothschild, Daniel G Goldstein, and Jake M Hofman. 2023b. Math Education with Large Language Models: Peril or Promise? Available at SSRN 4641653 (2023).
- Kumar et al. (2024) Harsh Kumar, Jonathan Vincentius, Ewan Jordan, and Ashton Anderson. 2024. Human Creativity in the Age of LLMs: Randomized Experiments on Divergent and Convergent Thinking. arXiv preprint arXiv:2410.03703 (2024).
- Lai et al. (2021) Vivian Lai, Chacha Chen, Q Vera Liao, Alison Smith-Renner, and Chenhao Tan. 2021. Towards a science of human-ai decision making: a survey of empirical studies. arXiv preprint arXiv:2112.11471 (2021).
- Lee and See (2004) John D Lee and Katrina A See. 2004. Trust in automation: Designing for appropriate reliance. Human factors 46, 1 (2004), 50–80.
- Liang et al. (2024) Weixin Liang, Zachary Izzo, Yaohui Zhang, Haley Lepp, Hancheng Cao, Xuandong Zhao, Lingjiao Chen, Haotian Ye, Sheng Liu, Zhi Huang, et al. 2024. Monitoring ai-modified content at scale: A case study on the impact of chatgpt on ai conference peer reviews. arXiv preprint arXiv:2403.07183 (2024).
- Liao and Varshney (2021) Q Vera Liao and Kush R Varshney. 2021. Human-centered explainable ai (xai): From algorithms to user experiences. arXiv preprint arXiv:2110.10790 (2021).
- Liao and Vaughan (2023) Q Vera Liao and Jennifer Wortman Vaughan. 2023. Ai transparency in the age of LLMs: A human-centered research roadmap. arXiv preprint arXiv:2306.01941 (2023), 5368–5393.
- Liao et al. (2022) Q Vera Liao, Yunfeng Zhang, Ronny Luss, Finale Doshi-Velez, and Amit Dhurandhar. 2022. Connecting algorithmic research and usage contexts: a perspective of contextualized evaluation for explainable AI. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 10. 147–159.
- Lin et al. (2022) Stephanie Lin, Jacob Hilton, and Owain Evans. 2022. Teaching models to express their uncertainty in words. arXiv preprint arXiv:2205.14334 (2022).
- Lo (2023) Chung Kwan Lo. 2023. What is the impact of ChatGPT on education? A rapid review of the literature. Education Sciences 13, 4 (2023), 410.
- Lu et al. (2023) Zhuoran Lu, Zhuoyan Li, Chun-Wei Chiang, and Ming Yin. 2023. Strategic Adversarial Attacks in AI-assisted Decision Making to Reduce Human Trust and Reliance.. In IJCAI. 3020–3028.
- Lu and Yin (2021) Zhuoran Lu and Ming Yin. 2021. Human reliance on machine learning models when performance feedback is limited: Heuristics and risks. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–16.
- Ma et al. (2024b) Qianou Ma, Hua Shen, Kenneth Koedinger, and Sherry Tongshuang Wu. 2024b. How to Teach Programming in the AI Era? Using LLMs as a Teachable Agent for Debugging. In International Conference on Artificial Intelligence in Education. Springer, 265–279.
- Ma et al. (2024a) Shuai Ma, Qiaoyi Chen, Xinru Wang, Chengbo Zheng, Zhenhui Peng, Ming Yin, and Xiaojuan Ma. 2024a. Towards human-ai deliberation: Design and evaluation of llm-empowered deliberative ai for ai-assisted decision-making. arXiv preprint arXiv:2403.16812 (2024).
- Ma et al. (2023) Shuai Ma, Ying Lei, Xinru Wang, Chengbo Zheng, Chuhan Shi, Ming Yin, and Xiaojuan Ma. 2023. Who should i trust: Ai or myself? leveraging human and ai correctness likelihood to promote appropriate trust in ai-assisted decision-making. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–19.
- Miller (2019) Tim Miller. 2019. Explanation in artificial intelligence: Insights from the social sciences. Artificial intelligence 267 (2019), 1–38.
- Misuraca et al. (2015) Raffaella Misuraca, Palmira Faraci, Amelia Gangemi, Floriana A Carmeci, and Silvana Miceli. 2015. The Decision Making Tendency Inventory: A new measure to assess maximizing, satisficing, and minimizing. Personality and Individual Differences 85 (2015), 111–116.
- Nam et al. (2024) Daye Nam, Andrew Macvean, Vincent Hellendoorn, Bogdan Vasilescu, and Brad Myers. 2024. Using an LLM to help with code understanding. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering. 1–13.
- Nguyen et al. (2018) An T Nguyen, Aditya Kharosekar, Saumyaa Krishnan, Siddhesh Krishnan, Elizabeth Tate, Byron C Wallace, and Matthew Lease. 2018. Believe it or not: designing a human-ai partnership for mixed-initiative fact-checking. In Proceedings of the 31st annual ACM symposium on user interface software and technology. 189–199.
- Nourani et al. (2021) Mahsan Nourani, Chiradeep Roy, Jeremy E Block, Donald R Honeycutt, Tahrima Rahman, Eric Ragan, and Vibhav Gogate. 2021. Anchoring bias affects mental model formation and user reliance in explainable AI systems. In Proceedings of the 26th International Conference on Intelligent User Interfaces. 340–350.
- Poursabzi-Sangdeh et al. (2021) Forough Poursabzi-Sangdeh, Daniel G Goldstein, Jake M Hofman, Jennifer Wortman Wortman Vaughan, and Hanna Wallach. 2021. Manipulating and measuring model interpretability. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–52.
- Reingold et al. (2024) Omer Reingold, Judy Hanwen Shen, and Aditi Talati. 2024. Dissenting Explanations: Leveraging Disagreement to Reduce Model Overreliance. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 21537–21544.
- Salikutluk et al. (2023) Vildan Salikutluk, Dorothea Koert, and Frank Jäkel. 2023. Interacting with large language models: A case study on AI-aided brainstorming for guesstimation problems. In HHAI 2023: Augmenting Human Intellect. IOS Press, 153–167.
- Schemmer et al. (2022) Max Schemmer, Patrick Hemmer, Niklas Kühl, Carina Benz, and Gerhard Satzger. 2022. Should I follow AI-based advice? Measuring appropriate reliance in human-AI decision-making. arXiv preprint arXiv:2204.06916 (2022).
- Schemmer et al. (2023) Max Schemmer, Niklas Kuehl, Carina Benz, Andrea Bartos, and Gerhard Satzger. 2023. Appropriate reliance on AI advice: Conceptualization and the effect of explanations. In Proceedings of the 28th International Conference on Intelligent User Interfaces. 410–422.
- Schoeffer et al. (2024) Jakob Schoeffer, Maria De-Arteaga, and Niklas Kuehl. 2024. Explanations, Fairness, and Appropriate Reliance in Human-AI Decision-Making. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–18.
- Sharma et al. (2024) Nikhil Sharma, Q Vera Liao, and Ziang Xiao. 2024. Generative Echo Chamber? Effect of LLM-Powered Search Systems on Diverse Information Seeking. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–17.
- Shen et al. (2024) Hua Shen, Tiffany Knearem, Reshmi Ghosh, Kenan Alkiek, Kundan Krishna, Yachuan Liu, Ziqiao Ma, Savvas Petridis, Yi-Hao Peng, Li Qiwei, et al. 2024. Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions. arXiv preprint arXiv:2406.09264 (2024).
- Si et al. (2023) Chenglei Si, Navita Goyal, Sherry Tongshuang Wu, Chen Zhao, Shi Feng, Hal Daumé III, and Jordan Boyd-Graber. 2023. Large Language Models Help Humans Verify Truthfulness–Except When They Are Convincingly Wrong. arXiv preprint arXiv:2310.12558 (2023).
- Sohrawardi et al. (2024) Saniat Javid Sohrawardi, Y Kelly Wu, Andrea Hickerson, and Matthew Wright. 2024. Dungeons & Deepfakes: Using scenario-based role-play to study journalists’ behavior towards using AI-based verification tools for video content. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–17.
- Spatharioti et al. (2023) Sofia Eleni Spatharioti, David M Rothschild, Daniel G Goldstein, and Jake M Hofman. 2023. Comparing traditional and llm-based search for consumer choice: A randomized experiment. arXiv preprint arXiv:2307.03744 (2023).
- Steyvers et al. (2024) Mark Steyvers, Heliodoro Tejeda, Aakriti Kumar, Catarina Belem, Sheer Karny, Xinyue Hu, Lukas Mayer, and Padhraic Smyth. 2024. The calibration gap between model and human confidence in large language models. arXiv preprint arXiv:2401.13835 (2024).
- Tankelevitch et al. (2024) Lev Tankelevitch, Viktor Kewenig, Auste Simkute, Ava Elizabeth Scott, Advait Sarkar, Abigail Sellen, and Sean Rintel. 2024. The metacognitive demands and opportunities of generative AI. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–24.
- Vafa et al. (2024) Keyon Vafa, Ashesh Rambachan, and Sendhil Mullainathan. 2024. Do Large Language Models Perform the Way People Expect? Measuring the Human Generalization Function. arXiv preprint arXiv:2406.01382 (2024).
- Vasconcelos et al. (2023a) Helena Vasconcelos, Gagan Bansal, Adam Fourney, Q Vera Liao, and Jennifer Wortman Vaughan. 2023a. Generation probabilities are not enough: Exploring the effectiveness of uncertainty highlighting in AI-powered code completions. arXiv preprint arXiv:2302.07248 (2023).
- Vasconcelos et al. (2023b) Helena Vasconcelos, Matthew Jörke, Madeleine Grunde-McLaughlin, Tobias Gerstenberg, Michael S Bernstein, and Ranjay Krishna. 2023b. Explanations can reduce overreliance on ai systems during decision-making. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (2023), 1–38.
- Virk et al. (2024) Yuvraj Virk, Premkumar Devanbu, and Toufique Ahmed. 2024. Enhancing Trust in LLM-Generated Code Summaries with Calibrated Confidence Scores. arXiv preprint arXiv:2404.19318 (2024).
- Vodrahalli et al. (2022) Kailas Vodrahalli, Tobias Gerstenberg, and James Y Zou. 2022. Uncalibrated models can improve human-ai collaboration. Advances in Neural Information Processing Systems 35 (2022), 4004–4016.
- Wang et al. (2022) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2022. Self-consistency improves chain of thought reasoning in language models. arXiv preprint arXiv:2203.11171 (2022).
- Wang and Yin (2021) Xinru Wang and Ming Yin. 2021. Are explanations helpful? A comparative study of the effects of explanations in ai-assisted decision-making. In Proceedings of the 26th International Conference on Intelligent User Interfaces. 318–328.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35 (2022), 24824–24837.
- Wolfe and Mitra (2024) Robert Wolfe and Tanushree Mitra. 2024. The Impact and Opportunities of Generative AI in Fact-Checking. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. 1531–1543.
- Xu et al. (2023) Ruiyun Xu, Yue Feng, and Hailiang Chen. 2023. Chatgpt vs. google: A comparative study of search performance and user experience. arXiv preprint arXiv:2307.01135 (2023).
- Yang et al. (2024) Blair Yang, Fuyang Cui, Keiran Paster, Jimmy Ba, Pashootan Vaezipoor, Silviu Pitis, and Michael R Zhang. 2024. Report Cards: Qualitative Evaluation of Language Models Using Natural Language Summaries. arXiv preprint arXiv:2409.00844 (2024).
- Yang et al. (2023) Rui Yang, Ting Fang Tan, Wei Lu, Arun James Thirunavukarasu, Daniel Shu Wei Ting, and Nan Liu. 2023. Large language models in health care: Development, applications, and challenges. Health Care Science 2, 4 (2023), 255–263.
- Zhai et al. (2024) Chunpeng Zhai, Santoso Wibowo, and Lily D Li. 2024. The effects of over-reliance on AI dialogue systems on students’ cognitive abilities: a systematic review. Smart Learning Environments 11, 1 (2024), 28.
- Zhang (2023) Peter Zhang. 2023. Taking advice from chatgpt. arXiv preprint arXiv:2305.11888 (2023).
- Zhang et al. (2023) Yue Zhang, Yafu Li, Leyang Cui, Deng Cai, Lemao Liu, Tingchen Fu, Xinting Huang, Enbo Zhao, Yu Zhang, Yulong Chen, et al. 2023. Siren’s song in the AI ocean: a survey on hallucination in large language models. arXiv preprint arXiv:2309.01219 (2023).
- Zhang et al. (2020) Yunfeng Zhang, Q Vera Liao, and Rachel KE Bellamy. 2020. Effect of confidence and explanation on accuracy and trust calibration in AI-assisted decision making. In Proceedings of the 2020 conference on fairness, accountability, and transparency. 295–305.
- Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
- Zhou et al. (2024) Kaitlyn Zhou, Jena D Hwang, Xiang Ren, Nouha Dziri, Dan Jurafsky, and Maarten Sap. 2024. Rel-AI: An Interaction-Centered Approach To Measuring Human-LM Reliance. arXiv preprint arXiv:2407.07950 (2024).
Appendix A Task Question Examples
Figures A.1 and A.2 show screenshots LSAT Logical Reasoning tasks questions. Participants answer independently in Stage 1, and then re-answer the same question with randomly sampled LLM advice in Stage 2.
Figures A.3 and A.4 show screenshots of Numerical Estimation task questions, with the same two-stage design.
Appendix B Control and Intervention Advice Examples
Appendix C Levels of Advice in Numerical Estimation
The five levels of LLM advice provided for the two questions in the Numerical Estimation task are illustrated in Figure C.7, alongside the distribution of the participants’ estimates before and after receiving the advice.
Appendix D Confidence Analysis for Other WoA Thresholds
The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.8 and D.9.
Appendix E Effect of Prior LLM Experience
As an additional analysis, we examine whether participants’ baseline frequencies of using LLM tools affected their behaviours and perceptions. The participants reported their prior experience with LLMs in the survey as never tried, almost never but have tried, monthly, weekly, and daily, which we map into ordinal values from 0 to 4. When controlling for condition, a partial Spearman correlation reveals no significant correlations between LLM usage and reliance on the advice for either task, LSAT ( for a binary reliance variable) or Numerical Estimation ( for the continuous WoA variable). Nor are there correlations with the changes in confidence for LSAT () and Numerical Estimation (). However, there is significant positive association between frequency of usage and all three categories of post-task perception: confidence (), accuracy (), and helpfulness (). This suggests that while people who were more familiar and open to using LLM tools generally hold higher view of ChatGPT’s advice (across all conditions), it did not impact actual rates of reliance in the task.

### v2

@@
 Our findings indicate that while interventions reduce over-reliance, they generally fail to improve appropriate reliance.
-Furthermore, people became more confident after making incorrect reliance decisions in certain contexts, demonstrating poor calibration.
+Furthermore, people became more confident after making wrong reliance decisions in certain contexts, demonstrating poor calibration.
 Based on our findings, we discuss implications for designing effective reliance interventions in human-LLM collaboration.
+Visual representation of four conditions.
+Each condition is represented by a text bubble coming from the right, with a text box on the left side.
+From left to right, the four conditions are Baseline, Reliance Disclaimer, Uncertainty Highlighting, and Implicit Answer.
+The textbox for Baseline contains “The answer is A” followed by lines representing unspecified text.
+The textbox for Reliance Disclaimer contains “The answer is A” followed by lines representing unspecified text, followed by “Remember to verify the information!” The textbox for Uncertainty Highlighting contains “The answer is A” followed by lines representing unspecified text, parts of which are highlighted in light pink or dark pink.
+The textbox for Baseline contains lines representing unspecified text.
 1. Introduction
@@
 LLMs encode remarkably vast amounts of knowledge; can communicate with users in clear, natural language; and have the potential to improve both the quality and efficiency of users’ decision-making (Tankelevitch et al., 2024).
-However, people can over-rely on LLM advice (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) or, in contrast, under-rely on it if they mistrust their LLM partner (Zhang, 2023).
-LLMs are prone to hallucination and error, yet often present their outputs with impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023).
-In contrast, algorithm aversion—the tendency to distrust algorithmic agents—can lead users to ignore helpful recommendations due to general distrust towards the technology (Dietvorst et al., 2015; Burton et al., 2020).
-Conscious effort must therefore be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024).
+While LLMs are prone to hallucination and error, they can still mislead users into a false sense of trust with their impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023).
+On the other hand, algorithm aversion can cause users to also ignore helpful recommendations due to a general distrust towards technology (Dietvorst et al., 2015; Burton et al., 2020).
+Consequently, both over-reliance (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) and under-reliance present significant yet contrasting challenges in human-LLM collaboration (Zhang, 2023).
+Conscious effort must be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024).
 Neither over-reliance nor under-reliance is ideal; instead, the focus should be to achieve appropriate reliance (Schemmer et al., 2022; Ma et al., 2023; He et al., 2023).
-Research in the tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021).
-In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in human-LLM collaboration.
-Due to the risk of hallucinations, reliance and calibration interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024).
-Reliance interventions that involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024), are being actively developed and evaluated.
-Since interventions are evolving rapidly, many of these methods still lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022).
-Furthermore, people’s subjective perceptions—such as their self-confidence—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration.
+Due to the risk of hallucinations, reliance interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance as well (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024).
+In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in question-answering tasks.
+The tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021).
+Reliance and calibration interventions for LLMs can involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024).
+Since interventions are being developed and evolving at a rapid pace, many of these methods lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022).
+Furthermore, people’s subjective perceptions—such as their self-confidence when receiving advice from LLMs—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration.
 Finally, the impact of different intervention types on reliance and perceptions remains unclear, as existing studies do not benchmark multiple interventions.
@@
 Factors that influence reliance include the expertise of the user in the task, their cognitive engagement in the task, and the information presented by the AI (Chen et al., 2023; He et al., 2023; Gajos and Mamykina, 2022).
-On the individual level, people may also be averse or appreciative towards algorithms, perhaps through pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023).
+On the individual level, people may also be averse or appreciative towards algorithms, perhaps due to pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023).
 A number of strategies have been developed to augment users’ engagement and to provide them with more information of the AI decision, such as through providing explanations (Reingold et al., 2024; Vasconcelos et al., 2023b; Zhang et al., 2020; Schoeffer et al., 2024).
@@
 Recent research has investigated the role of LLMs in human-LLM collaborations for applications such as writing (Liang et al., 2024), communication (Fu et al., 2024), coding (Kazemitabaar et al., 2023; Nam et al., 2024), education (Lo, 2023; Kumar et al., 2023b; Bo et al., 2024), fact-checking (Wolfe and Mitra, 2024; Kim et al., 2024a; Si et al., 2023), information search (Spatharioti et al., 2023; Salikutluk et al., 2023; Sharma et al., 2024), and other specialized domains (Choudhury and Shamszare, 2023; Cheong et al., 2024; Yang et al., 2023).
-In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations (Zhang, 2023).
+In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations.
 They can also offer antagonistic advice to induce the user into considering alternative options (Chiang et al., 2024; Ma et al., 2024a; Cai et al., 2024).
@@
 In technical AI research, the quality of LLM responses has been improved through methods such as chain-of-thought with intermediate reasoning steps (Wei et al., 2022), debating amongst multiple LLMs (Khan et al., 2024; Kim et al., 2024a), and enforcing self-consistency in outputs (Wang et al., 2022; Becker and Soatto, 2024).
-In many cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022).
+In ideal cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022).
 However, the question of how to induce appropriate reliance is still open.
@@
 3. Methodology
-We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We selected GPT-4o (https://openai.com/index/hello-gpt-4o/) as it is the most state-of-art LLM at the time of experimentation.
+Two figures labelled a) diagram depicting the experimental procedure as explained in Methods, and b) sample of LSAT task question with the corresponding advice that each of the four conditions saw.
+We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We used gpt-4o-2024-08-06 as it was the most state-of-art GPT model at the time of experimentation.
 in two question-answering tasks.
-To answer our research questions, we will compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the baseline LLM advice in the Control condition.
+To answer our research questions, we compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the effect of the baseline Control condition.
 In this section, we outline our choices of tasks, procedure, measures, and conditions.
@@
 Based on preliminary results from piloting multiple-choice questions from several domains, we found that logical reasoning questions best balanced difficulty with accessibility.
-We selected LSAT logical reasoning questions as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’, for the first task.
-Each LSAT question has five options, of which exactly one is correct.
+We selected LSAT logical reasoning questions222Law School Admission Test® (LSAT) questions created by Law School Admission Council (LSAC).
+as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’.
+Each LSAT question has five options, of which exactly one is correct (determined based on LSAT answer keys).
 We generated both correct and incorrect LLM advice for each question for use during the experiment.
-The questions and some sample LLM advice can be found in Figures A.1-A.2 in Appendix A.
-Correct answers were determined using LSAT answer keys.
-While LSAT questions are highly detailed, they are designed to test for aptitude in law school rather than knowledge of the legal domain.
+A sample question and corresponding LLM advice for all four conditions can be found in Figure 2(b), and screenshots of what participants saw can be found in Figures 1(b) in Appendix A.
+While LSAT questions can be difficult to comprehend, they are designed to test for aptitude in law school rather than knowledge of the legal domain.
 As such, they do not contain undefined jargon, nor do they require domain knowledge (satisfying D1).
-In addition to reasoning, LSAT Logical Reasoning questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2).
+In addition to reasoning, LSAT questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2).
 Furthermore, we reworded a question so that entering it into a search engine would not result in the correct answer (satisfying D3).
@@
 Numerical Estimation.
-While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a numerical estimation task adds a valuable new dimension to this research space.
-Among the various forms of numerical estimation, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar.
+While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a Numerical Estimation (NE) task adds a valuable new dimension to this research space.
+Among the various forms of NE tasks, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar.
 We posit that this is a gamified task that is familiar to most people and does not require advanced mathematics (satisfying D1).
@@
 One set featured a smaller jar containing 403 chickpeas, while the other contained a larger jar with 856 jellybeans.
-The pictures and sample LLM advice can be found in Figures A.3-A.4 in Appendix A.
+The pictures and sample LLM advice can be found in Figures A.2-A.3 in Appendix A.
 Unlike in multiple-choice tasks, where advice can be objectively classified as good or bad, the goodness of the advice in the Numerical Estimation setting depends on the participant’s initial estimate.
@@
 In the experiment, each participant was randomly assigned to one of four conditions and then answered two questions from the LSAT task and two questions from the Numerical Estimation task, randomized so that the LSAT phase and Numerical Estimation phase were equally likely to appear first.
-The orders of questions within each category were also randomized for each participant.
+The orders of questions within each category were also randomized.
 Participants received pre-generated advice from GPT-4o.
-This advice was randomly sampled for each question, meaning participants could receive any mix of good or bad advice.
+This advice was randomly sampled for each question, meaning participants could receive any mix of ‘good’ or ‘bad’ advice.
 After completing the tasks, participants completed a short post-survey.
-See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice from all four conditions.
+See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice for all four conditions for an LSAT question (for the advice provided in an Numerical Estimation question, see Figure B.2).
 To mimic the stakes and level of engagement typically involved in real-world decision-making scenarios where people might consult AI chatbots, we incorporated two key design elements into the study:
@@
 Participants earn rewards for correct answers, with two opportunities for bonus pay: one before, and one after receiving LLM-generated advice.
-In the LSAT questions, participants received a bonus if they selected the correct answer, while for the numerical estimation task, participants earned a bonus if they guessed within 10% of the true count.
+In the LSAT questions, participants received a bonus if they selected the correct answer, while for the Numerical Estimation questions, participants earned a bonus if they guessed within 10% of the true count.
 3.3. Measures
@@
 See Figure 3 for an illustration of the reliance decisions pathways.
+Flowchart demonstrating reliance pathways.
+There are three categories of elements: Initial Outcome, Action, and Reliance Behaviour.
+The diagram depicts the categories of Negative Self-Reliance (NS), Positive LLM Reliance (PL), Positive Self-Reliance (PS), and Negative LLM Reliance (NL) as described in the paper.
+Regardless of Action, “LLM advice is the exact same as Initial Answer” flows to “No reliance”.
 What does it mean to rely on LLM advice?
 In multiple-choice questions, we follow previous work by defining reliance as changing one’s answer to the LLM’s suggestion.
-In estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as .
+In numerical estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as — in short, this means that the participant was influenced by the LLM’s advice to adjust their estimate.
 We further define the reliance outcomes as falling into one of four categories:
@@
 Positive Self-Reliance (PS): the participant does not follow the LLM’s advice, and doing so would have worsened or failed to improve their performance.
-Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance.
-Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance.
+Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance. This is also referred to as over-reliance.
+Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance. This is also referred to as under-reliance.
 PL and PS can be grouped together as Correct Reliance Behaviours, while NL and NS are Wrong Reliance Behaviours.
 NL represents over-reliance on LLM advice (the participant took the advice when it worsens or fails to improve improve their performance) and NS represents under-reliance (the participant didn’t take the advice but should have).
-We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extend to other real-world tasks.
+We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extended to other real-world tasks.
 In contrast, the LSAT task lacks relative scoring across options, treating all incorrect answers as equally wrong; so relying on incorrect advice counts as over-reliance.
@@
 From these behaviour classifications, we compute the following appropriate reliance metrics:
-Relative LLM Reliance (RLR) = Rate of appropriate reliance on correct LLM advice.
-Relative Self-Reliance (RSR) = Rate of appropriate self-reliance when the LLM advice is wrong.
-Appropriate Reliance Ratio (ARR) = Ratio of appropriate reliance on correct advice to over-reliance on wrong advice, where:
-ARR ¡ 1: Higher rate of over-reliance than positive reliance.
+Relative LLM Reliance (RLR) = Rate of correctly relying on the LLM when the LLM advice is helpful.
+Relative Self-Reliance (RSR) = Rate of correctly relying on oneself when the LLM advice is harmful.
+Appropriate Reliance Ratio (ARR) = Ratio of correct reliance on helpful advice to wrong reliance on harmful advice, where:
+ARR ¡ 1: Higher rate of over-reliance than positive reliance, signifying inappropriate reliance behaviour.
 ARR = 1: Random chance; LLM reliance rates are the same regardless of LLM correctness.
-ARR ¿ 1: Higher rate of positive reliance than over-reliance.
+ARR ¿ 1: Higher rate of positive reliance than over-reliance, signifying appropriate reliance behaviour.
 We compute the metrics for each condition across participants, then calculate the effect size as the difference in outcomes between the control and each intervention.
@@
 Confidence and Subjective Perceptions.
-For each question, we collect the participant’s confidence in the answer they gave prior to receiving LLM advice (Initial Confidence), and their confidence in the answer they gave after (Final Confidence) receiving the LLM-generated advice.
-Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence from the first to second stage.
-Well-calibrated confidence should result in lower confidence scores when participants make incorrect reliance decisions.
+For each question, we collect the participant’s confidence in their answer prior to receiving LLM advice (Initial Confidence), and their confidence in their answer after receiving the LLM advice (Final Confidence) .
+Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence as a result of receiving LLM advice.
+Ideally, well-calibrated confidence should result in lower confidence scores when participants make wrong reliance decisions.
 We compare the change in confidence between correct and wrong reliance outcomes for each of the conditions, where statistical differences are tested with Mann-Whitney U-test for non-normal distributions with significance threshold .
 At the end of the survey, participants provide additional ratings on a bipolar Likert scale of 1 (disagree) to 7 (agree) for the following subjective perception statements:
-ChatGPT’s advice made me more confident.222This subjective confidence measure is separate from the question-level confidence measure (the actual target of RQ2).
+ChatGPT’s advice made me more confident333Note that this measure of post-task confidence is different from the question-level confidence (the target of RQ2) described previously..
 ChatGPT’s advice seems accurate.
@@
 Task Performance.
-Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative task performance.
+Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative performance.
 However, we explore the initial error from participants independently answering questions, and compare this with their updated error after re-answering questions with LLM advice.
@@
 Time of Engagement.
-To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on answering each stage of question-answering.
-From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, and the time taken to answer the initial question (Relative Advice Time).
+To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on each stage of question-answering.
+From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, to the time taken to answer the initial question (Relative Time).
+A larger value potentially signifies higher engagement in the advice, which may correlate with the complexity of the intervention paired with the advice.
 3.4. Experimental Conditions
@@
 To generate the baseline advice, we input the questions into GPT-4o through OpenAI’s API and record the responses directly.
-Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to an explanation supporting the answer.
+Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to a 2-4 sentence long explanation supporting the answer.
 Reliance Disclaimer.
-The following static disclaimer statement is appended to all advice generated in Control: “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”.
-This text was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully.
+We append the following static disclaimer statement, “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”, while keeping the advice generated in Control the same.
+This disclaimer was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully.
 For consistency and to emulate real-world chatbot disclaimers like ChatGPT, we apply the same disclaimer to all advice.
@@
 Uncertainty Highlighting.
-When generating the Control advice, we record the log probability of each token.
-Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a).
+When generating the Control advice, we also record the log probability of each token.
 To visually augment the baseline advice, we highlight the tokens with log probability under -1 in stronger red and indicate them to be uncertain, and the tokens under -0.5 in lighter pink and indicate them to be slightly uncertain.
@@
 However, they were reminded of the definition of the highlighting colors in every survey page with the LLM advice.
+Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a).
 Implicit Answer.
-This intervention semantically augments the content of the advice to imply the answer without writing it directly, requiring more engagement to extract the answer that the advice supports.
-For the Logical Reasoning questions, we append an instruction to the prompt to never explicitly state the answer’s text, and to use paraphrasing to generate an argument in support of the answer.
+This intervention semantically augments the content of the advice to avoid writing the answer directly, requiring more engagement to extract which answer the advice supports.
+For the Logical Reasoning questions, we instruct GPT-4o to never explicitly state the answer, and to use paraphrasing to generate an argument in support of the answer.
 For the Numerical Estimation questions, we manually removed the final calculation step from the advice generated in Control (as it was difficult to achieve this through prompting alone).
-In this way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
+This way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
 4. Results
-We pre-registered our experimental design and analysis plan online333https://aspredicted.org/8BD_WWN, then ran the experiment.
-Here, we report the analyses described in Section 3 separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
+We pre-registered our experimental design and analysis plan online 444https://aspredicted.org/8BD_WWN before conducting the experiments.
+Here, we report the results of the analyses described in Section 3, separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
 4.1. Participants
@@
 Furthermore, they are well-distributed in their usage of AI chatbots, with 18.9% citing daily use, 33.5% citing weekly use, 17.9% citing monthly use, and the remaining 29.7% citing rare or no usage.
+Performance outcomes for LSAT questions.
+For each intervention condition, there is a “transition matrix” showing the proportion of people who answered wrongly initially then answered wrongly with LLM advice, wrongly initially then answered correctly with LLM advice, correctly initially then answered wrongly with LLM advice, correctly initially then answered correctly with LLM advice.
+Two figures labelled (a) and (b) depict percent error distributions and WoA distributions in Numerical Estimation.
+Subfigure (a) depicts the distribution of Stage 1 and Stage 2 percent error via four graphs: one for each intervention condition.
+Subfigure (b) depicts the distribution of weight of advice via four graphs: one for each intervention condition.
 4.2. Task Performance
@@
 In each question, if the human-LLM team relied on the team member that got the correct answer, then the Best Possible Performance would reach 62.1% accuracy (not 100%, because sometimes both the human and the LLM answered incorrectly).
-The gap between optimal and actual performance represents the performance lost to inappropriate reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
+The gap between optimal and actual performance represents the performance lost to wrong reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
 Numerical Estimation.
@@
 While unrelated to task performance, we further compute how much the participants relied on the advice for additional insights on their behaviour.
-The weight-of-advice taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions.
+The weight-of-advice (WoA) taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions.
 The distribution of WoA tends to be bimodal, with participants in Control, Reliance Disclaimer, and Uncertainty Highlighting predominantly favouring adoption of the LLM advice.
@@
 By the Mann-Whitney U test, Reliance Disclaimer’s WoA () and Implicit Answer’s WoA () are significantly smaller than Control’s WoA.
+Three reliance figures labelled (a), (b), and (c).
+All of them have Relative LLM Reliance (RLR) on the x-axis, and Relative Self-Reliance (RSR) on the y-axis.
+Figure (a) depicts reference reliance outcomes.
+Figure (b) depicts reliance outcomes for LSAT questions.
+Figure (c) depicts reliance outcomes for Numerical Estimation questions.
 4.3. Reliance Behaviour
-Following the methods of previous work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases.
+Following the method of prior work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases (see the third pathway in Figure 3).
 For Numerical Estimation, we do not use a strict definition of agreement, as it is unlikely for a participant’s initial estimate to exactly match the LLM advice.
@@
 For example, a participant may perceive a guess of 700 to be “pretty much the same” as an advice of 750.
-We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.444We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small.
+We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.555We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small.
 To address this, we include an additional 50 padding around the advice.
@@
 To answer RQ1, we compute the Relative LLM Reliance (RLR), Relative Self-Reliance (RSR), and Appropriate Reliance Ratio (ARR) across all data instances collected per task in each condition.
-For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is bad.
-ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM (less over-reliance and less under-reliance, or at least a positive trade-off such as much less over-reliance and only a small increase in under-reliance).
+For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is harmful.
+ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM.
 In terms of reliance outcomes, over-reliance maps to low RSR (i.e.
@@
 Table 1 displays all the metrics calculated for each condition, separately for the LSAT and Numerical Estimation tasks.
-For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control.
+For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control666This study followed a preregistered analysis plan using as the significance threshold.
+While no multiple-hypothesis corrections were applied (as per the preregistration), we acknowledge that adjusting for multiple tests could influence the significance level..
 Control | Reliance Disclaimer | Uncertainty Highlighting | Implicit Answer | ||
@@
 ARR | 1.45 | 1.91, p=.2 | 1.34, p=.4 | 1.03, p=.1 |
+Depicts RLR, RSR, and ARR for each intervention category and each question type–showing Numerical estimation reliance metrics with WoA thresholds of 0.01, 0.5, and 0.99.
+The interventions with statistically significant reliance metric differences generally show increases in RSR or decreases in RLR.
+Only Reliance Disclaimer has a statistically significant increase in ARR.
 LSAT.
 The visual representations of the metrics and their significances for LSAT are plotted in Figure 6(b), where Control is plotted in gray and we compare the interventions against it.
-We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance (i.e.
-it is higher and further to the right than Control).
-Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance, but at the cost of shifting reliance away from the LLM when following it would be appropriate (although this decrease is not statistically significant)—these interventions tend to reduce LLM reliance overall, both when the advice is bad and when it is good.
-Both interventions have higher ARR values than Control, but the difference is not significant at our power level.
-These results indicate that a static, persistent disclaimer message can induce appropriate reliance more effectively than more specialized methods.
+We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance ().
+Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance ( and , respectively), but at the cost of shifting overall reliance away from the LLM.
+Despite so, the reduction of RLR – signifying correct reliance on LLM – is not significant.
+Both interventions have higher ARR values than Control, but the difference is again not significant at our power level.
+These results indicate that a static, persistent disclaimer message may induce appropriate reliance more effectively than more specialized methods.
 Numerical Estimation.
@@
 Values and significance for RLR, RSS, and ARR are expressed at these three WoA threshold levels in both Table 1 and Figure 6(c).
-Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance.
-Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 and 0.5).
-At the lowest WoA threshold, this also increases under-reliance—which is undesirable.
+Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance through increasing ARR.
+Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 with and 0.5 at ).
+At the lowest WoA threshold, this also increases under-reliance through significantly reducing RLR ()—which is undesirable.
 This could be due to being forced to perform simple math calculations to see the advised number—which people may not do if they are satisfied with their own answer, or believe that extracting the advice requires more effort than the reward is worth.
-Uncertainty Highlighting also causes significant decrease in RLR at high WoA thresholds, indicating that people are hesitant to fully use the advice.
+Uncertainty Highlighting also causes significant decrease in RLR at a high WoA threshold of 0.99 (), indicating that people are hesitant to fully use the advice.
 This corroborates Figure 5(b), which shows that Uncertainty Highlighting and Reliance Disclaimer both have reduced density of WoA values close to 1 compared to Control.
@@
 The discrepancy in results between our two tasks also warns that results of interventions may not generalize beyond the specific tasks they are tested on, emphasizing the need for thorough validation across diverse tasks before broader deployment.
+LSAT confidence changes by reliance behaviour.
+Two double bar graphs labelled (a) and (b).
+Both depict average changes in confidence levels when answering LSAT questions.
+Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
+Average change in confidence is positive except for Correct Reliance Behaviour under the Highlighting condition.
+Only Highlighting has statistically different changes in confidence for right vs.
+wrong reliance behaviours, with positive confidence changes after wrong reliance behaviour and barely decreasing average confidence after right reliance behaviour.
+Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
+Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
+Both are grouped by condition.
+The average change in confidence is always positive.
+Only Disclaimer shows a statistically significant difference, with average change in confidence after a correct answer being significantly higher after appropriate reliance than after overreliance.
+NE confidence changes by reliance behaviour, with a WoA threshold of 0.5.
+Two double bar graphs labelled (a) and (b).
+Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.5.
+Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
+Average change in confidence is always positive.
+Only Implicit Answer has statistically different changes in confidence for right vs.
+wrong reliance behaviours, with confidence changes close to 0 for wrong reliance behaviour and increases in confidence for right reliance behaviour.
+Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
+Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
+Both are grouped by condition.
+The average change in confidence is always positive.
+There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
 4.4. Confidence Calibration and Perceptions
-To address RQ2, we examine participant confidence in their pre-advice answer and their post-advice answer for each question, along with several post-experiment Likert scale ratings of their subjective perceptions.
-We compare the changes in confidence between cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (over-reliance).
+To address RQ2, we compare participants’ confidence in their pre-advice answer and their post-advice answer (Change in Confidence) at the question level, along with several post-experiment Likert scale ratings of their subjective perceptions.
+We compare cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (cases where participants actively relied on the LLM).
 Unless otherwise indicated, significance is calculated within each condition using the Mann-Whitney U test with as the significance threshold.
@@
 LSAT.
-Figure 7(a) shows participants’ LSAT question confidence changes for correct and wrong reliance behaviours.
-The most salient takeaway is that on-aggregate across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions ().
+Figure 7(a) shows participants’ Change in Confidence in the LSAT task for correct and wrong reliance behaviours.
+The most salient takeaway is that on-aggregate, across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions ().
 Within each condition, Uncertainty Highlighting is the only one where this observed effect is significant (), which indicates poor confidence calibration since it is undesirable for people to be more confident in wrong decisions.
@@
 Here, the confidence scores tend to increase more as people trusted the LLMs enough to use them.
-We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is significantly lower when the participant over-relied than when they relied appropriately ().
-Though not significant, it is interesting to observe that the opposite trend can be demonstrated in the other interventions, where confidence increases more in cases of over-reliance.
+We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is lower when the participant over-relied than when they relied appropriately ().
+Though not significant, it is interesting to observe that the opposite trend can be demonstrated in some of the other interventions, where confidence increases more in cases of over-reliance (NL).
 This indicates that reliance interventions that operate by moderating users’ confidence in the LLM may inadvertently miscalibrate confidence with respect to appropriate reliance.
@@
 For this task, we repeat the confidence analyses and show them in Figure 8.
-For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.8-D.9).
-Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes.
-However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs, which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
+For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.1-D.2).
+Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes (Figure 8(a)).
+However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs (Figure 8(b)), which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
 Overall Post-Perception.
@@
 An alternative approach to moderate this effect may be to highlight highly certain tokens in green in addition to uncertain tokens in red.
-In addition, we find that the frequency of LLM usage correlates positively with all three subjective perception measures, but not with reliance behaviours – see Appendix E for details.
+In addition, we find that the frequency of baseline LLM usage correlates positively with all three subjective perception measures – see Appendix E for the details of this analysis.
 The main implication here is that uncertainty calibration methods can have the undesirable outcome of inducing negative perceptions of accuracy and usefulness of the LLM, which increases overall algorithm aversion against LLM advice, rather than appropriate reliance.
+A quadruple bar graph depicting post-experiment perceptions.
+Each colour represents a different intervention condition (Control, Disclaimer, Highlighting, Implicit Answer).
+All three post-experiment perceptions are depicted.
+Two bar graphs labelled (a), and (b) showing average Time Ratios for each question type.
+The y-axes are both labelled Time Ratio.
+Both x-axes have Control, Disclaimer, Highlighting, and Implicit Answer as categories.
+Bar graph (a) shows the average Time Ratios for LSAT questions, for which Highlighting and Implicit Answer are close to 0.6 and Control and Disclaimer are close to 0.4.
+Bar graph (b) shows the average Time Ratios for Numerical Estimation questions, for which Implicit Answer is the highest at around 1.6, Disclaimer and Highlighting are close to 1.0, and Control is close to 0.8.
 4.5. Time of Engagement
@@
 Our results suggest that the added friction of calculating the answer in Implicit Answer caused people to disproportionately discard the LLM advice (not rely at all), while the confidence calibration in Uncertainty Highlighting caused people to hesitate to fully follow advice (rely, but not fully).
-The increase in advice under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b).
+The increase in under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b).
 This suggests that for the LSAT task, static Reliance Disclaimers may have added just enough friction to dissuade habitual over-reliance.
@@
 We examine how participants’ confidence in their answers change after seeing the LLM’s advice.
-Ideally, even with incorrect reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice.
+Ideally, even with wrong reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice.
 However, we largely observe the opposite or inconclusive effects, where wrong reliance results in higher gains in confidence.
@@
 Since participants knew that they were participating in an experiment, they may have inferred that they were being evaluated on LLM reliance and become more cautious about their reliance behaviours.
-Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to “hedge”; to maximize their chances of earning some bonus payment by choosing different answers at each stage.
+Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to ‘hedge’; to maximize their chances of earning some bonus payment by choosing different answers at each stage.
 Our two-stage question answering set-up also emulates a specific type of decision-making where users are already engaged in the task.
@@
 The distribution of advice quality was held constant regardless of the LLM’s true tendencies when responding to the question to prevent question difficulty from being a confound.
-We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude555https://claude.ai/, Copilot666https://copilot.microsoft.com/, Gemini777https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour.
+We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude777https://claude.ai/, Copilot888https://copilot.microsoft.com/, Gemini999https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour.
 We also did not explore varying the system prompts to steer the LLM into embodying different personas.
@@
 Appendix A Task Question Examples
-Figures A.1 and A.2 show screenshots LSAT Logical Reasoning tasks questions.
+Figure 1(b) shows an example of the LSAT Logical Reasoning task question.
 Participants answer independently in Stage 1, and then re-answer the same question with randomly sampled LLM advice in Stage 2.
-Figures A.3 and A.4 show screenshots of Numerical Estimation task questions, with the same two-stage design.
+The “Road” LSAT Logical Reasoning question is depicted via two figures labelled (a) and (b).
+Screenshot (a) depicts Stage 1.
+A multiple-choice question’s text presents a paragraph making an argument, then asks “Which of the following, if true, most strengthens the argument in the paragraph above?” Options A-E are shown below.
+Screenshot (b) depicts Stage 2.
+Text at the top of this screenshot says “Now you have access to ChatGPT’s advice.
+Will you change your answer?
+Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same question from above is repeated.
+Below we see sample advice from ChatGPT followed by “Please enter what you believe is the correct answer”.
+Below that, the multiple choice options are displayed again.
+Figures A.2 and A.3 show screenshots of Numerical Estimation task questions, with the same two-stage design.
+The Chickpea Numerical Estimation question is depicted via two Qualtrics screenshots labelled (a) and (b).
+Screenshot (a) depicts Stage 1.
+Question text asks “How many chickpeas are in this glass jar?” Two pictures are displayed: a bottom-up view and a side view of the same jar of chickpeas.
+Below we see “Please enter a number representing the number of chickpeas:” followed by a text box.
+Screenshot (b) depicts Stage 2.
+Text at the top of this screenshot says “Now you have access to ChatGPT’s advice.
+Will you change your answer?
+Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same two pictures of the chickpea jar are shown.
+Below, we see the participant’s Stage 1 answer.
+Below this, we see sample advice from ChatGPT.
+Below that, we see “Please enter a number representing the number of chickpeas:” followed by a text box.
+The Jellybean Numerical Estimation question is depicted via two Qualtrics screenshots labelled (a) and (b).
+Screenshot (a) depicts Stage 1.
+Question text asks “How many jellybeans are in this glass jar?” Two pictures are displayed: a bottom-up view and a side view of the same jar of jellybeans.
+Below we see “Please enter a number representing the number of jellybeans:” followed by a text box.
+Screenshot (b) depicts Stage 2.
+Text at the top of this screenshot says “Now you have access to ChatGPT’s advice.
+Will you change your answer?
+Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same two pictures of the jellybean jar are shown.
+Below we see the participant’s Stage 1 answer.
+Below this, we see sample advice from ChatGPT.
+Below that, we see “Please enter a number representing the number of jellybeans:” followed by a text box.
 Appendix B Control and Intervention Advice Examples
+Figures 1(d) and B.2 show sample ChatGPT advice for all four conditions for a LSAT Logical Reasoning question and a Numerical Estimation, respectively.
+Four Qualtrics screenshots labelled (a), (b), (c), and (d) with sample advice for the sample LSAT question for each condition.
+Screenshot (a) depicts Control condition advice, which recommends option B with reasoning.
+Screenshot (b) depicts Reliance Disclaimer advice, which is identical to Control condition advice but followed with bolded text saying “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.” Screenshot (c) depicts Uncertainty Highlighting advice, which is identical to Control condition advice but with some text that is highlighted pink and dark pink.
+Screenshot (d) depicts Implicit Answer advice, which has reasoning for why Option B is correct–without explicitly referring to Option B.
+Four Qualtrics screenshots labelled (a), (b), (c), and (d) with sample advice for the sample NE question for each condition.
+Screenshot (a) depicts Control condition advice, which has a 4-step answer.
+Step 1 estimates chickpea dimensions.
+Step 2 makes assumptions about jar dimensions and calculates its volume.
+Step 3 estimates chickpea volume.
+Step 4 divides jar volume by chickpea volume to get an estimate for total chickpeas, concluding that there are approximately 887 chickpeas.
+Screenshot (b) depicts Reliance Disclaimer advice, which is identical to Control condition advice but followed with bolded text saying “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.” Screenshot (c) depicts Uncertainty Highlighting advice, which is identical to Control condition advice but with some text that is highlighted pink and dark pink.
+Screenshot (d) depicts Implicit Answer advice, which is identical to Control condition advice but with the final step explained rather than shown.
 Appendix C Levels of Advice in Numerical Estimation
 The five levels of LLM advice provided for the two questions in the Numerical Estimation task are illustrated in Figure C.1, alongside the distribution of the participants’ estimates before and after receiving the advice.
+Two figures labelled (a) and (b) depicting the distribution of participant estimations and LLM advice levels for the Numerical Estimation Task, separately for all four conditions.
+Subfigure (a) is the chickpea estimation question and subfigure (b) is the jelly bean estimation question.
 Appendix D Confidence Analysis for Other WoA Thresholds
-The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.8 and D.9.
+The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.1 and D.2.
+NE confidence changes by reliance behaviour, with a WoA threshold of 0.01.
+Two double bar graphs labelled (a) and (b).
+Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.99.
+Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
+Average change in confidence is always positive.
+For each category, there is no statistically significant difference between average change in confidence after wrong vs.
+correct reliance behaviour.
+Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
+Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
+Both are grouped by condition.
+The average change in confidence is always positive.
+There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
+NE confidence changes by reliance behaviour, with a WoA threshold of 0.99.
+Two double bar graphs labelled (a) and (b).
+Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.99.
+Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
+Average change in confidence is always positive.
+Only Implicit Answer has statistically different changes in confidence for right vs.
+wrong reliance behaviours, with confidence changes close to 0 for wrong reliance behaviour and increases in confidence for right reliance behaviour.
+Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
+Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
+Both are grouped by condition.
+The average change in confidence is always positive.
+There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
 Appendix E Effect of Prior LLM Experience
@@
 Nor are there correlations with the changes in confidence for LSAT () and Numerical Estimation ().
-However, there is significant positive association between frequency of usage and all three categories of post-task perception: confidence (), accuracy (), and helpfulness ().
+However, there is significant positive association between frequency of usage and all three categories of post-task perceptions of the LLM: Confidence (), Accuracy (), and Helpfulness ().
 This suggests that while people who were more familiar and open to using LLM tools generally hold higher view of ChatGPT’s advice (across all conditions), it did not impact actual rates of reliance in the task.

### v3

@@
 Our findings indicate that while interventions reduce over-reliance, they generally fail to improve appropriate reliance.
-Furthermore, people became more confident after making wrong reliance decisions in certain contexts, demonstrating poor calibration.
+Furthermore, people became more confident after making incorrect reliance decisions in certain contexts, demonstrating poor calibration.
 Based on our findings, we discuss implications for designing effective reliance interventions in human-LLM collaboration.
-Visual representation of four conditions.
-Each condition is represented by a text bubble coming from the right, with a text box on the left side.
-From left to right, the four conditions are Baseline, Reliance Disclaimer, Uncertainty Highlighting, and Implicit Answer.
-The textbox for Baseline contains “The answer is A” followed by lines representing unspecified text.
-The textbox for Reliance Disclaimer contains “The answer is A” followed by lines representing unspecified text, followed by “Remember to verify the information!” The textbox for Uncertainty Highlighting contains “The answer is A” followed by lines representing unspecified text, parts of which are highlighted in light pink or dark pink.
-The textbox for Baseline contains lines representing unspecified text.
 1. Introduction
@@
 LLMs encode remarkably vast amounts of knowledge; can communicate with users in clear, natural language; and have the potential to improve both the quality and efficiency of users’ decision-making (Tankelevitch et al., 2024).
-While LLMs are prone to hallucination and error, they can still mislead users into a false sense of trust with their impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023).
-On the other hand, algorithm aversion can cause users to also ignore helpful recommendations due to a general distrust towards technology (Dietvorst et al., 2015; Burton et al., 2020).
-Consequently, both over-reliance (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) and under-reliance present significant yet contrasting challenges in human-LLM collaboration (Zhang, 2023).
-Conscious effort must be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024).
+However, people can over-rely on LLM advice (Spatharioti et al., 2023; Kabir et al., 2024; Si et al., 2023) or, in contrast, under-rely on it if they mistrust their LLM partner (Zhang, 2023).
+LLMs are prone to hallucination and error, yet often present their outputs with impressive detail and confidence (Zhang et al., 2023; Chen and Shu, 2023).
+In contrast, algorithm aversion—the tendency to distrust algorithmic agents—can lead users to ignore helpful recommendations due to general distrust towards the technology (Dietvorst et al., 2015; Burton et al., 2020).
+Conscious effort must therefore be placed on helping users calibrate appropriate levels of trust and reliance on LLM advice (Zhang et al., 2020; Zhou et al., 2024).
 Neither over-reliance nor under-reliance is ideal; instead, the focus should be to achieve appropriate reliance (Schemmer et al., 2022; Ma et al., 2023; He et al., 2023).
-Due to the risk of hallucinations, reliance interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance as well (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024).
-In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in question-answering tasks.
-The tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021).
-Reliance and calibration interventions for LLMs can involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024).
-Since interventions are being developed and evolving at a rapid pace, many of these methods lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022).
-Furthermore, people’s subjective perceptions—such as their self-confidence when receiving advice from LLMs—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration.
+Research in the tangential field of human-AI collaboration has explored methods like explanations and calibration techniques to improve joint decision-making by enhancing AI transparency and verifiability, though results often show mixed outcomes and trade-offs (Vasconcelos et al., 2023b; Liao and Varshney, 2021; Miller, 2019; Schoeffer et al., 2024; Bansal et al., 2021).
+In this work, we adapt rigorous measures of appropriate reliance from prior works in human-AI collaboration (Schemmer et al., 2023) to benchmark the impact of several LLM reliance interventions in human-LLM collaboration.
+Due to the risk of hallucinations, reliance and calibration interventions in LLMs are often motivated by curbing over-relying on wrongful information (Kim et al., 2024b; Si et al., 2023; Spatharioti et al., 2023), however recent attention has also shifted to optimizing for appropriate reliance (Zhou et al., 2024; Zhang, 2023; Collins et al., 2024; Do et al., 2024).
+Reliance interventions that involve communicating the capabilities and confidence of an LLM to its users (Spatharioti et al., 2023), or introducing various types of frictions that the user must bypass to receive the advice (Collins et al., 2024), are being actively developed and evaluated.
+Since interventions are evolving rapidly, many of these methods still lack comprehensive evaluations of appropriate reliance in real users, especially those focusing on technical contributions (Lin et al., 2022; Becker and Soatto, 2024; Bowman et al., 2022).
+Furthermore, people’s subjective perceptions—such as their self-confidence—can have a complex relationship with their reliance behaviour (Kim et al., 2024b; Cao and Huang, 2022; Lu and Yin, 2021; Si et al., 2023), which warrants further exploration.
 Finally, the impact of different intervention types on reliance and perceptions remains unclear, as existing studies do not benchmark multiple interventions.
@@
 Factors that influence reliance include the expertise of the user in the task, their cognitive engagement in the task, and the information presented by the AI (Chen et al., 2023; He et al., 2023; Gajos and Mamykina, 2022).
-On the individual level, people may also be averse or appreciative towards algorithms, perhaps due to pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023).
+On the individual level, people may also be averse or appreciative towards algorithms, perhaps through pre-conceptions or judgments formed after seeing the AI’s performance (Kim et al., 2023; Dietvorst et al., 2015; Lu et al., 2023; Nourani et al., 2021; Ma et al., 2023).
 A number of strategies have been developed to augment users’ engagement and to provide them with more information of the AI decision, such as through providing explanations (Reingold et al., 2024; Vasconcelos et al., 2023b; Zhang et al., 2020; Schoeffer et al., 2024).
@@
 Recent research has investigated the role of LLMs in human-LLM collaborations for applications such as writing (Liang et al., 2024), communication (Fu et al., 2024), coding (Kazemitabaar et al., 2023; Nam et al., 2024), education (Lo, 2023; Kumar et al., 2023b; Bo et al., 2024), fact-checking (Wolfe and Mitra, 2024; Kim et al., 2024a; Si et al., 2023), information search (Spatharioti et al., 2023; Salikutluk et al., 2023; Sharma et al., 2024), and other specialized domains (Choudhury and Shamszare, 2023; Cheong et al., 2024; Yang et al., 2023).
-In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations.
+In these tasks, the LLM can provide guidance and information for the user to assist in their decision-making, learning, and creations (Zhang, 2023).
 They can also offer antagonistic advice to induce the user into considering alternative options (Chiang et al., 2024; Ma et al., 2024a; Cai et al., 2024).
@@
 In technical AI research, the quality of LLM responses has been improved through methods such as chain-of-thought with intermediate reasoning steps (Wei et al., 2022), debating amongst multiple LLMs (Khan et al., 2024; Kim et al., 2024a), and enforcing self-consistency in outputs (Wang et al., 2022; Becker and Soatto, 2024).
-In ideal cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022).
+In many cases, the human-LLM team can reach better performance as a team than individually, tapping into the strength of their combined knowledge and skill sets (Bowman et al., 2022).
 However, the question of how to induce appropriate reliance is still open.
@@
 3. Methodology
-Two figures labelled a) diagram depicting the experimental procedure as explained in Methods, and b) sample of LSAT task question with the corresponding advice that each of the four conditions saw.
-We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We used gpt-4o-2024-08-06 as it was the most state-of-art GPT model at the time of experimentation.
+We designed a large-scale, randomized, and incentivized experiment to evaluate the efficacy of three reliance interventions applied to the outputs of GPT-4o111We selected GPT-4o (https://openai.com/index/hello-gpt-4o/) as it is the most state-of-art LLM at the time of experimentation.
 in two question-answering tasks.
-To answer our research questions, we compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the effect of the baseline Control condition.
+To answer our research questions, we will compare the effect of the interventions on performance, reliance behaviour, confidence calibration, and time of engagement against the baseline LLM advice in the Control condition.
 In this section, we outline our choices of tasks, procedure, measures, and conditions.
@@
 Based on preliminary results from piloting multiple-choice questions from several domains, we found that logical reasoning questions best balanced difficulty with accessibility.
-We selected LSAT logical reasoning questions222Law School Admission Test® (LSAT) questions created by Law School Admission Council (LSAC).
-as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’.
-Each LSAT question has five options, of which exactly one is correct (determined based on LSAT answer keys).
+We selected LSAT logical reasoning questions as the domain for the multiple-choice task and adapted two multiple-choice LSAT logical reasoning questions, both labeled ‘difficult’, for the first task.
+Each LSAT question has five options, of which exactly one is correct.
 We generated both correct and incorrect LLM advice for each question for use during the experiment.
-A sample question and corresponding LLM advice for all four conditions can be found in Figure 2(b), and screenshots of what participants saw can be found in Figures 1(b) in Appendix A.
-While LSAT questions can be difficult to comprehend, they are designed to test for aptitude in law school rather than knowledge of the legal domain.
+The questions and some sample LLM advice can be found in Figures A.1-A.2 in Appendix A.
+Correct answers were determined using LSAT answer keys.
+While LSAT questions are highly detailed, they are designed to test for aptitude in law school rather than knowledge of the legal domain.
 As such, they do not contain undefined jargon, nor do they require domain knowledge (satisfying D1).
-In addition to reasoning, LSAT questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2).
+In addition to reasoning, LSAT Logical Reasoning questions also test analytical and reading comprehension skills, making them non-trivial to answer (satisfying D2).
 Furthermore, we reworded a question so that entering it into a search engine would not result in the correct answer (satisfying D3).
@@
 Numerical Estimation.
-While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a Numerical Estimation (NE) task adds a valuable new dimension to this research space.
-Among the various forms of NE tasks, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar.
+While most previous studies in human-AI collaboration focus on multiple-choice or binary outcome questions, we believe that introducing a numerical estimation task adds a valuable new dimension to this research space.
+Among the various forms of numerical estimation, such as predicting values in regression tasks, we opted for an image-based format based on a classic carnival game: estimating the quantity of beans in a glass jar.
 We posit that this is a gamified task that is familiar to most people and does not require advanced mathematics (satisfying D1).
@@
 One set featured a smaller jar containing 403 chickpeas, while the other contained a larger jar with 856 jellybeans.
-The pictures and sample LLM advice can be found in Figures A.2-A.3 in Appendix A.
+The pictures and sample LLM advice can be found in Figures A.3-A.4 in Appendix A.
 Unlike in multiple-choice tasks, where advice can be objectively classified as good or bad, the goodness of the advice in the numerical estimation setting depends on the participant’s initial estimate.
@@
 In the experiment, each participant was randomly assigned to one of four conditions and then answered two questions from the LSAT task and two questions from the Numerical Estimation task, randomized so that the LSAT phase and Numerical Estimation phase were equally likely to appear first.
-The orders of questions within each category were also randomized.
+The orders of questions within each category were also randomized for each participant.
 Participants received pre-generated advice from GPT-4o.
-This advice was randomly sampled for each question, meaning participants could receive any mix of ‘good’ or ‘bad’ advice.
+This advice was randomly sampled for each question, meaning participants could receive any mix of good or bad advice.
 After completing the tasks, participants completed a short post-survey.
-See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice for all four conditions for an LSAT question (for the advice provided in an Numerical Estimation question, see Figure B.2).
+See Figure 2(a) for the full diagram of the experimental flow and Figure 2(b) for sample advice from all four conditions.
 To mimic the stakes and level of engagement typically involved in real-world decision-making scenarios where people might consult AI chatbots, we incorporated two key design elements into the study:
@@
 Participants earn rewards for correct answers, with two opportunities for bonus pay: one before, and one after receiving LLM-generated advice.
-In the LSAT questions, participants received a bonus if they selected the correct answer, while for the Numerical Estimation questions, participants earned a bonus if they guessed within 10% of the true count.
+In the LSAT questions, participants received a bonus if they selected the correct answer, while for the numerical estimation task, participants earned a bonus if they guessed within 10% of the true count.
 3.3. Measures
@@
 See Figure 3 for an illustration of the reliance decisions pathways.
-Flowchart demonstrating reliance pathways.
-There are three categories of elements: Initial Outcome, Action, and Reliance Behaviour.
-The diagram depicts the categories of Negative Self-Reliance (NS), Positive LLM Reliance (PL), Positive Self-Reliance (PS), and Negative LLM Reliance (NL) as described in the paper.
-Regardless of Action, “LLM advice is the exact same as Initial Answer” flows to “No reliance”.
 What does it mean to rely on LLM advice?
 In multiple-choice questions, we follow previous work by defining reliance as changing one’s answer to the LLM’s suggestion.
-In numerical estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as — in short, this means that the participant was influenced by the LLM’s advice to adjust their estimate.
+In estimation questions, we define reliance as an answer with positive weight-of-advice (WoA), where WoA is calculated as .
 We further define the reliance outcomes as falling into one of four categories:
@@
 Positive Self-Reliance (PS): the participant does not follow the LLM’s advice, and doing so would have worsened or failed to improve their performance.
-Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance. This is also referred to as over-reliance.
-Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance. This is also referred to as under-reliance.
+Negative LLM Reliance (NL): the participant changes their original answer to follow the LLM’s advice, but worsens or fails to improve their performance.
+Negative Self-Reliance (NS): the participant does not follow the LLM’s advice, but doing so would have improved their performance.
 PL and PS can be grouped together as Correct Reliance Behaviours, while NL and NS are Wrong Reliance Behaviours.
 NL represents over-reliance on LLM advice (the participant took the advice when it worsens or fails to improve improve their performance) and NS represents under-reliance (the participant didn’t take the advice but should have).
-We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extended to other real-world tasks.
+We note that PL is possible even if the LLM advice is incorrect, as long as there is a notion of relative error—such as switching to a less wrong answer suggested by the LLM, which applies to the Numerical Estimation task and can extend to other real-world tasks.
 In contrast, the LSAT task lacks relative scoring across options, treating all incorrect answers as equally wrong; so relying on incorrect advice counts as over-reliance.
@@
 From these behaviour classifications, we compute the following appropriate reliance metrics:
-Relative LLM Reliance (RLR) = Rate of correctly relying on the LLM when the LLM advice is helpful.
-Relative Self-Reliance (RSR) = Rate of correctly relying on oneself when the LLM advice is harmful.
-Appropriate Reliance Ratio (ARR) = Ratio of correct reliance on helpful advice to wrong reliance on harmful advice, where:
-ARR ¡ 1: Higher rate of over-reliance than positive reliance, signifying inappropriate reliance behaviour.
+Relative LLM Reliance (RLR) = Rate of appropriate reliance on correct LLM advice.
+Relative Self-Reliance (RSR) = Rate of appropriate self-reliance when the LLM advice is wrong.
+Appropriate Reliance Ratio (ARR) = Ratio of appropriate reliance on correct advice to over-reliance on wrong advice, where:
+ARR ¡ 1: Higher rate of over-reliance than positive reliance.
 ARR = 1: Random chance; LLM reliance rates are the same regardless of LLM correctness.
-ARR ¿ 1: Higher rate of positive reliance than over-reliance, signifying appropriate reliance behaviour.
+ARR ¿ 1: Higher rate of positive reliance than over-reliance.
 We compute the metrics for each condition across participants, then calculate the effect size as the difference in outcomes between the control and each intervention.
@@
 Confidence and Subjective Perceptions.
-For each question, we collect the participant’s confidence in their answer prior to receiving LLM advice (Initial Confidence), and their confidence in their answer after receiving the LLM advice (Final Confidence) .
-Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence as a result of receiving LLM advice.
-Ideally, well-calibrated confidence should result in lower confidence scores when participants make wrong reliance decisions.
+For each question, we collect the participant’s confidence in the answer they gave prior to receiving LLM advice (Initial Confidence), and their confidence in the answer they gave after (Final Confidence) receiving the LLM-generated advice.
+Confidence is measured using a unipolar Likert scale from 1 (highly uncertain) to 5 (highly certain), which allows us to calculate the Change in Confidence from the first to second stage.
+Well-calibrated confidence should result in lower confidence scores when participants make incorrect reliance decisions.
 We compare the change in confidence between correct and wrong reliance outcomes for each of the conditions, where statistical differences are tested with Mann-Whitney U-test for non-normal distributions with significance threshold .
 At the end of the survey, participants provide additional ratings on a bipolar Likert scale of 1 (disagree) to 7 (agree) for the following subjective perception statements:
-ChatGPT’s advice made me more confident333Note that this measure of post-task confidence is different from the question-level confidence (the target of RQ2) described previously..
+ChatGPT’s advice made me more confident.222This subjective confidence measure is separate from the question-level confidence measure (the actual target of RQ2).
 ChatGPT’s advice seems accurate.
@@
 Task Performance.
-Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative performance.
+Since our experiment is not designed to emulate the realistic performance of LLMs in tasks, we do not particularly emphasize the human-LLM collaborative task performance.
 However, we explore the initial error from participants independently answering questions, and compare this with their updated error after re-answering questions with LLM advice.
@@
 Time of Engagement.
-To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on each stage of question-answering.
-From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, to the time taken to answer the initial question (Relative Time).
-A larger value potentially signifies higher engagement in the advice, which may correlate with the complexity of the intervention paired with the advice.
+To understand participant engagement with the LLM’s advice, we measure the amount of time that participants spend on answering each stage of question-answering.
+From these two time measurements, we calculate the relative time as the ratio between time taken to engage with the advice, and the time taken to answer the initial question (Relative Advice Time).
 3.4. Experimental Conditions
@@
 To generate the baseline advice, we input the questions into GPT-4o through OpenAI’s API and record the responses directly.
-Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to a 2-4 sentence long explanation supporting the answer.
+Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to an explanation supporting the answer.
 Reliance Disclaimer.
-We append the following static disclaimer statement, “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”, while keeping the advice generated in Control the same.
-This disclaimer was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully.
+The following static disclaimer statement is appended to all advice generated in Control: “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”.
+This text was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully.
 For consistency and to emulate real-world chatbot disclaimers like ChatGPT, we apply the same disclaimer to all advice.
@@
 Uncertainty Highlighting.
-When generating the Control advice, we also record the log probability of each token.
+When generating the Control advice, we record the log probability of each token.
+Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a).
 To visually augment the baseline advice, we highlight the tokens with log probability under -1 in stronger red and indicate them to be uncertain, and the tokens under -0.5 in lighter pink and indicate them to be slightly uncertain.
@@
 However, they were reminded of the definition of the highlighting colors in every survey page with the LLM advice.
-Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a).
 Implicit Answer.
-This intervention semantically augments the content of the advice to avoid writing the answer directly, requiring more engagement to extract which answer the advice supports.
-For the Logical Reasoning questions, we instruct GPT-4o to never explicitly state the answer, and to use paraphrasing to generate an argument in support of the answer.
+This intervention semantically augments the content of the advice to imply the answer without writing it directly, requiring more engagement to extract the answer that the advice supports.
+For the Logical Reasoning questions, we append an instruction to the prompt to never explicitly state the answer’s text, and to use paraphrasing to generate an argument in support of the answer.
 For the Numerical Estimation questions, we manually removed the final calculation step from the advice generated in Control (as it was difficult to achieve this through prompting alone).
-This way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
+In this way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
 4. Results
-We pre-registered our experimental design and analysis plan online 444https://aspredicted.org/8BD_WWN before conducting the experiments.
-Here, we report the results of the analyses described in Section 3, separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
+We pre-registered our experimental design and analysis plan online333https://aspredicted.org/8BD_WWN, then ran the experiment.
+Here, we report the analyses described in Section 3 separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
 4.1. Participants
@@
 Furthermore, they are well-distributed in their usage of AI chatbots, with 18.9% citing daily use, 33.5% citing weekly use, 17.9% citing monthly use, and the remaining 29.7% citing rare or no usage.
-Performance outcomes for LSAT questions.
-For each intervention condition, there is a “transition matrix” showing the proportion of people who answered wrongly initially then answered wrongly with LLM advice, wrongly initially then answered correctly with LLM advice, correctly initially then answered wrongly with LLM advice, correctly initially then answered correctly with LLM advice.
-Two figures labelled (a) and (b) depict percent error distributions and WoA distributions in Numerical Estimation.
-Subfigure (a) depicts the distribution of Stage 1 and Stage 2 percent error via four graphs: one for each intervention condition.
-Subfigure (b) depicts the distribution of weight of advice via four graphs: one for each intervention condition.
 4.2. Task Performance
@@
 In each question, if the human-LLM team relied on the team member that got the correct answer, then the Best Possible Performance would reach 62.1% accuracy (not 100%, because sometimes both the human and the LLM answered incorrectly).
-The gap between optimal and actual performance represents the performance lost to wrong reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
+The gap between optimal and actual performance represents the performance lost to inappropriate reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
 Numerical Estimation.
@@
 While unrelated to task performance, we further compute how much the participants relied on the advice for additional insights on their behaviour.
-The weight-of-advice (WoA) taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions.
+The weight-of-advice taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions.
 The distribution of WoA tends to be bimodal, with participants in Control, Reliance Disclaimer, and Uncertainty Highlighting predominantly favouring adoption of the LLM advice.
@@
 By the Mann-Whitney U test, Reliance Disclaimer’s WoA () and Implicit Answer’s WoA () are significantly smaller than Control’s WoA.
-Three reliance figures labelled (a), (b), and (c).
-All of them have Relative LLM Reliance (RLR) on the x-axis, and Relative Self-Reliance (RSR) on the y-axis.
-Figure (a) depicts reference reliance outcomes.
-Figure (b) depicts reliance outcomes for LSAT questions.
-Figure (c) depicts reliance outcomes for Numerical Estimation questions.
 4.3. Reliance Behaviour
-Following the method of prior work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases (see the third pathway in Figure 3).
+Following the methods of previous work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases.
 For Numerical Estimation, we do not use a strict definition of agreement, as it is unlikely for a participant’s initial estimate to exactly match the LLM advice.
@@
 For example, a participant may perceive a guess of 700 to be “pretty much the same” as an advice of 750.
-We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.555We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small.
+We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.444We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small.
 To address this, we include an additional 50 padding around the advice.
@@
 To answer RQ1, we compute the Relative LLM Reliance (RLR), Relative Self-Reliance (RSR), and Appropriate Reliance Ratio (ARR) across all data instances collected per task in each condition.
-For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is harmful.
-ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM.
+For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is bad.
+ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM (less over-reliance and less under-reliance, or at least a positive trade-off such as much less over-reliance and only a small increase in under-reliance).
 In terms of reliance outcomes, over-reliance maps to low RSR (i.e.
@@
 Table 1 displays all the metrics calculated for each condition, separately for the LSAT and Numerical Estimation tasks.
-For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control666This study followed a preregistered analysis plan using as the significance threshold.
-While no multiple-hypothesis corrections were applied (as per the preregistration), we acknowledge that adjusting for multiple tests could influence the significance level..
-Control | Reliance Disclaimer | Uncertainty Highlighting | Implicit Answer | ||
-LSAT | RLR | 0.66 | 0.71, p=.2 | 0.52, p=.1 | 0.58, p=.2 |
-RSR | 0.44 | 0.56, p=.05 | 0.62, p=.01 (+) | 0.63, p=.002 (+) | |
-ARR | 1.17 | 1.62, p=.04 (+) | 1.35, p=.2 | 1.54, p=.09 | |
-Numerical Estimation (WoA >0.01) | RLR | 0.91 | 0.91, p=.4 | 0.93, p=.3 | 0.67, p<.001 (–) |
-RSR | 0.15 | 0.22, p=.1 | 0.23, p=.1 | 0.43, p=.004 (+) | |
-ARR | 1.07 | 1.17, p=.2 | 1.20, p=.1 | 1.18, r=.2 | |
-Numerical Estimation (WoA >0.5) | RLR | 0.8 | 0.74, p=.2 | 0.8, p=.5 | 0.45, p=.17 |
-RSR | 0.39 | 0.47, p=.1 | 0.42, p=.3 | 0.55, p=.007 (+) | |
-ARR | 1.32 | 1.38, p=.4 | 1.37, p=.4 | 1.01, p=.1 | |
-Numerical Estimation (WoA >0.99) | RLR | 0.47 | 0.39, p=.2 | 0.3, p=.02 (–) | 0.34, p=.05 |
-RSR | 0.68 | 0.79, p=.02 (+) | 0.77, p=.05 | 0.67, p=.5 | |
+For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control.
+| Control | Reliance Disclaimer | Uncertainty Highlighting | Implicit Answer | |||||
+| LSAT | RLR | 0.66 | 0.71, p=.2 | 0.52, p=.1 | 0.58, p=.2 | |||
+| RSR | 0.44 | 0.56, p=.05 | 0.62, p=.01 (+) | 0.63, p=.002 (+) | ||||
+| ARR | 1.17 | 1.62, p=.04 (+) | 1.35, p=.2 | 1.54, p=.09 | ||||
+RLR | 0.91 | 0.91, p=.4 | 0.93, p=.3 | 0.67, p<.001 (–) | |||
+| RSR | 0.15 | 0.22, p=.1 | 0.23, p=.1 | 0.43, p=.004 (+) | ||||
+| ARR | 1.07 | 1.17, p=.2 | 1.20, p=.1 | 1.18, r=.2 | ||||
+RLR | 0.8 | 0.74, p=.2 | 0.8, p=.5 | 0.45, p=.17 | |||
+| RSR | 0.39 | 0.47, p=.1 | 0.42, p=.3 | 0.55, p=.007 (+) | ||||
+| ARR | 1.32 | 1.38, p=.4 | 1.37, p=.4 | 1.01, p=.1 | ||||
+RLR | 0.47 | 0.39, p=.2 | 0.3, p=.02 (–) | 0.34, p=.05 | |||
+| RSR | 0.68 | 0.79, p=.02 (+) | 0.77, p=.05 | 0.67, p=.5 | ||||
 | ARR | 1.45 | 1.91, p=.2 | 1.34, p=.4 | 1.03, p=.1 |
-Depicts RLR, RSR, and ARR for each intervention category and each question type–showing Numerical estimation reliance metrics with WoA thresholds of 0.01, 0.5, and 0.99.
-The interventions with statistically significant reliance metric differences generally show increases in RSR or decreases in RLR.
-Only Reliance Disclaimer has a statistically significant increase in ARR.
 LSAT.
 The visual representations of the metrics and their significances for LSAT are plotted in Figure 6(b), where Control is plotted in gray and we compare the interventions against it.
-We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance ().
-Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance ( and , respectively), but at the cost of shifting overall reliance away from the LLM.
-Despite so, the reduction of RLR – signifying correct reliance on LLM – is not significant.
-Both interventions have higher ARR values than Control, but the difference is again not significant at our power level.
-These results indicate that a static, persistent disclaimer message may induce appropriate reliance more effectively than more specialized methods.
+We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance (i.e.
+it is higher and further to the right than Control).
+Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance, but at the cost of shifting reliance away from the LLM when following it would be appropriate (although this decrease is not statistically significant)—these interventions tend to reduce LLM reliance overall, both when the advice is bad and when it is good.
+Both interventions have higher ARR values than Control, but the difference is not significant at our power level.
+These results indicate that a static, persistent disclaimer message can induce appropriate reliance more effectively than more specialized methods.
 Numerical Estimation.
@@
 Values and significance for RLR, RSS, and ARR are expressed at these three WoA threshold levels in both Table 1 and Figure 6(c).
-Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance through increasing ARR.
-Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 with and 0.5 at ).
-At the lowest WoA threshold, this also increases under-reliance through significantly reducing RLR ()—which is undesirable.
+Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance.
+Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 and 0.5).
+At the lowest WoA threshold, this also increases under-reliance—which is undesirable.
 This could be due to being forced to perform simple math calculations to see the advised number—which people may not do if they are satisfied with their own answer, or believe that extracting the advice requires more effort than the reward is worth.
-Uncertainty Highlighting also causes significant decrease in RLR at a high WoA threshold of 0.99 (), indicating that people are hesitant to fully use the advice.
+Uncertainty Highlighting also causes significant decrease in RLR at high WoA thresholds, indicating that people are hesitant to fully use the advice.
 This corroborates Figure 5(b), which shows that Uncertainty Highlighting and Reliance Disclaimer both have reduced density of WoA values close to 1 compared to Control.
@@
 The discrepancy in results between our two tasks also warns that results of interventions may not generalize beyond the specific tasks they are tested on, emphasizing the need for thorough validation across diverse tasks before broader deployment.
-LSAT confidence changes by reliance behaviour.
-Two double bar graphs labelled (a) and (b).
-Both depict average changes in confidence levels when answering LSAT questions.
-Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
-Average change in confidence is positive except for Correct Reliance Behaviour under the Highlighting condition.
-Only Highlighting has statistically different changes in confidence for right vs.
-wrong reliance behaviours, with positive confidence changes after wrong reliance behaviour and barely decreasing average confidence after right reliance behaviour.
-Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
-Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
-Both are grouped by condition.
-The average change in confidence is always positive.
-Only Disclaimer shows a statistically significant difference, with average change in confidence after a correct answer being significantly higher after appropriate reliance than after overreliance.
-NE confidence changes by reliance behaviour, with a WoA threshold of 0.5.
-Two double bar graphs labelled (a) and (b).
-Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.5.
-Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
-Average change in confidence is always positive.
-Only Implicit Answer has statistically different changes in confidence for right vs.
-wrong reliance behaviours, with confidence changes close to 0 for wrong reliance behaviour and increases in confidence for right reliance behaviour.
-Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
-Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
-Both are grouped by condition.
-The average change in confidence is always positive.
-There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
 4.4. Confidence Calibration and Perceptions
-To address RQ2, we compare participants’ confidence in their pre-advice answer and their post-advice answer (Change in Confidence) at the question level, along with several post-experiment Likert scale ratings of their subjective perceptions.
-We compare cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (cases where participants actively relied on the LLM).
+To address RQ2, we examine participant confidence in their pre-advice answer and their post-advice answer for each question, along with several post-experiment Likert scale ratings of their subjective perceptions.
+We compare the changes in confidence between cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (over-reliance).
 Unless otherwise indicated, significance is calculated within each condition using the Mann-Whitney U test with as the significance threshold.
@@
 LSAT.
-Figure 7(a) shows participants’ Change in Confidence in the LSAT task for correct and wrong reliance behaviours.
-The most salient takeaway is that on-aggregate, across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions ().
+Figure 7(a) shows participants’ LSAT question confidence changes for correct and wrong reliance behaviours.
+The most salient takeaway is that on-aggregate across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions ().
 Within each condition, Uncertainty Highlighting is the only one where this observed effect is significant (), which indicates poor confidence calibration since it is undesirable for people to be more confident in wrong decisions.
@@
 Here, the confidence scores tend to increase more as people trusted the LLMs enough to use them.
-We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is lower when the participant over-relied than when they relied appropriately ().
-Though not significant, it is interesting to observe that the opposite trend can be demonstrated in some of the other interventions, where confidence increases more in cases of over-reliance (NL).
+We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is significantly lower when the participant over-relied than when they relied appropriately ().
+Though not significant, it is interesting to observe that the opposite trend can be demonstrated in the other interventions, where confidence increases more in cases of over-reliance.
 This indicates that reliance interventions that operate by moderating users’ confidence in the LLM may inadvertently miscalibrate confidence with respect to appropriate reliance.
@@
 For this task, we repeat the confidence analyses and show them in Figure 8.
-For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.1-D.2).
-Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes (Figure 8(a)).
-However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs (Figure 8(b)), which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
+For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.8-D.9).
+Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes.
+However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs, which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
 Overall Post-Perception.
@@
 An alternative approach to moderate this effect may be to highlight highly certain tokens in green in addition to uncertain tokens in red.
-In addition, we find that the frequency of baseline LLM usage correlates positively with all three subjective perception measures – see Appendix E for the details of this analysis.
+In addition, we find that the frequency of LLM usage correlates positively with all three subjective perception measures, but not with reliance behaviours – see Appendix E for details.
 The main implication here is that uncertainty calibration methods can have the undesirable outcome of inducing negative perceptions of accuracy and usefulness of the LLM, which increases overall algorithm aversion against LLM advice, rather than appropriate reliance.
-A quadruple bar graph depicting post-experiment perceptions.
-Each colour represents a different intervention condition (Control, Disclaimer, Highlighting, Implicit Answer).
-All three post-experiment perceptions are depicted.
-Two bar graphs labelled (a), and (b) showing average Time Ratios for each question type.
-The y-axes are both labelled Time Ratio.
-Both x-axes have Control, Disclaimer, Highlighting, and Implicit Answer as categories.
-Bar graph (a) shows the average Time Ratios for LSAT questions, for which Highlighting and Implicit Answer are close to 0.6 and Control and Disclaimer are close to 0.4.
-Bar graph (b) shows the average Time Ratios for Numerical Estimation questions, for which Implicit Answer is the highest at around 1.6, Disclaimer and Highlighting are close to 1.0, and Control is close to 0.8.
 4.5. Time of Engagement
@@
 Our results suggest that the added friction of calculating the answer in Implicit Answer caused people to disproportionately discard the LLM advice (not rely at all), while the confidence calibration in Uncertainty Highlighting caused people to hesitate to fully follow advice (rely, but not fully).
-The increase in under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b).
+The increase in advice under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b).
 This suggests that for the LSAT task, static Reliance Disclaimers may have added just enough friction to dissuade habitual over-reliance.
@@
 We examine how participants’ confidence in their answers change after seeing the LLM’s advice.
-Ideally, even with wrong reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice.
+Ideally, even with incorrect reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice.
 However, we largely observe the opposite or inconclusive effects, where wrong reliance results in higher gains in confidence.
@@
 Since participants knew that they were participating in an experiment, they may have inferred that they were being evaluated on LLM reliance and become more cautious about their reliance behaviours.
-Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to ‘hedge’; to maximize their chances of earning some bonus payment by choosing different answers at each stage.
+Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to “hedge”; to maximize their chances of earning some bonus payment by choosing different answers at each stage.
 Our two-stage question answering set-up also emulates a specific type of decision-making where users are already engaged in the task.
@@
 The distribution of advice quality was held constant regardless of the LLM’s true tendencies when responding to the question to prevent question difficulty from being a confound.
-We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude777https://claude.ai/, Copilot888https://copilot.microsoft.com/, Gemini999https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour.
+We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude555https://claude.ai/, Copilot666https://copilot.microsoft.com/, Gemini777https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour.
 We also did not explore varying the system prompts to steer the LLM into embodying different personas.
@@
 Appendix A Task Question Examples
-Figure 1(b) shows an example of the LSAT Logical Reasoning task question.
+Figures A.1 and A.2 show screenshots LSAT Logical Reasoning tasks questions.
 Participants answer independently in Stage 1, and then re-answer the same question with randomly sampled LLM advice in Stage 2.
-The “Road” LSAT Logical Reasoning question is depicted via two figures labelled (a) and (b).
-Screenshot (a) depicts Stage 1.
-A multiple-choice question’s text presents a paragraph making an argument, then asks “Which of the following, if true, most strengthens the argument in the paragraph above?” Options A-E are shown below.
-Screenshot (b) depicts Stage 2.
-Text at the top of this screenshot says “Now you have access to ChatGPT’s advice.
-Will you change your answer?
-Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same question from above is repeated.
-Below we see sample advice from ChatGPT followed by “Please enter what you believe is the correct answer”.
-Below that, the multiple choice options are displayed again.
-Figures A.2 and A.3 show screenshots of Numerical Estimation task questions, with the same two-stage design.
-The Chickpea Numerical Estimation question is depicted via two Qualtrics screenshots labelled (a) and (b).
-Screenshot (a) depicts Stage 1.
-Question text asks “How many chickpeas are in this glass jar?” Two pictures are displayed: a bottom-up view and a side view of the same jar of chickpeas.
-Below we see “Please enter a number representing the number of chickpeas:” followed by a text box.
-Screenshot (b) depicts Stage 2.
-Text at the top of this screenshot says “Now you have access to ChatGPT’s advice.
-Will you change your answer?
-Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same two pictures of the chickpea jar are shown.
-Below, we see the participant’s Stage 1 answer.
-Below this, we see sample advice from ChatGPT.
-Below that, we see “Please enter a number representing the number of chickpeas:” followed by a text box.
-The Jellybean Numerical Estimation question is depicted via two Qualtrics screenshots labelled (a) and (b).
-Screenshot (a) depicts Stage 1.
-Question text asks “How many jellybeans are in this glass jar?” Two pictures are displayed: a bottom-up view and a side view of the same jar of jellybeans.
-Below we see “Please enter a number representing the number of jellybeans:” followed by a text box.
-Screenshot (b) depicts Stage 2.
-Text at the top of this screenshot says “Now you have access to ChatGPT’s advice.
-Will you change your answer?
-Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same two pictures of the jellybean jar are shown.
-Below we see the participant’s Stage 1 answer.
-Below this, we see sample advice from ChatGPT.
-Below that, we see “Please enter a number representing the number of jellybeans:” followed by a text box.
+Figures A.3 and A.4 show screenshots of Numerical Estimation task questions, with the same two-stage design.
 Appendix B Control and Intervention Advice Examples
-Figures 1(d) and B.2 show sample ChatGPT advice for all four conditions for a LSAT Logical Reasoning question and a Numerical Estimation, respectively.
-Four Qualtrics screenshots labelled (a), (b), (c), and (d) with sample advice for the sample LSAT question for each condition.
-Screenshot (a) depicts Control condition advice, which recommends option B with reasoning.
-Screenshot (b) depicts Reliance Disclaimer advice, which is identical to Control condition advice but followed with bolded text saying “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.” Screenshot (c) depicts Uncertainty Highlighting advice, which is identical to Control condition advice but with some text that is highlighted pink and dark pink.
-Screenshot (d) depicts Implicit Answer advice, which has reasoning for why Option B is correct–without explicitly referring to Option B.
-Four Qualtrics screenshots labelled (a), (b), (c), and (d) with sample advice for the sample NE question for each condition.
-Screenshot (a) depicts Control condition advice, which has a 4-step answer.
-Step 1 estimates chickpea dimensions.
-Step 2 makes assumptions about jar dimensions and calculates its volume.
-Step 3 estimates chickpea volume.
-Step 4 divides jar volume by chickpea volume to get an estimate for total chickpeas, concluding that there are approximately 887 chickpeas.
-Screenshot (b) depicts Reliance Disclaimer advice, which is identical to Control condition advice but followed with bolded text saying “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.” Screenshot (c) depicts Uncertainty Highlighting advice, which is identical to Control condition advice but with some text that is highlighted pink and dark pink.
-Screenshot (d) depicts Implicit Answer advice, which is identical to Control condition advice but with the final step explained rather than shown.
 Appendix C Levels of Advice in Numerical Estimation
 The five levels of LLM advice provided for the two questions in the Numerical Estimation task are illustrated in Figure C.7, alongside the distribution of the participants’ estimates before and after receiving the advice.
-Two figures labelled (a) and (b) depicting the distribution of participant estimations and LLM advice levels for the Numerical Estimation Task, separately for all four conditions.
-Subfigure (a) is the chickpea estimation question and subfigure (b) is the jelly bean estimation question.
 Appendix D Confidence Analysis for Other WoA Thresholds
-The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.1 and D.2.
-NE confidence changes by reliance behaviour, with a WoA threshold of 0.01.
-Two double bar graphs labelled (a) and (b).
-Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.99.
-Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
-Average change in confidence is always positive.
-For each category, there is no statistically significant difference between average change in confidence after wrong vs.
-correct reliance behaviour.
-Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
-Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
-Both are grouped by condition.
-The average change in confidence is always positive.
-There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
-NE confidence changes by reliance behaviour, with a WoA threshold of 0.99.
-Two double bar graphs labelled (a) and (b).
-Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.99.
-Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition.
-Average change in confidence is always positive.
-Only Implicit Answer has statistically different changes in confidence for right vs.
-wrong reliance behaviours, with confidence changes close to 0 for wrong reliance behaviour and increases in confidence for right reliance behaviour.
-Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2.
-Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect).
-Both are grouped by condition.
-The average change in confidence is always positive.
-There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
+The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.8 and D.9.
 Appendix E Effect of Prior LLM Experience
@@
 Nor are there correlations with the changes in confidence for LSAT () and Numerical Estimation ().
-However, there is significant positive association between frequency of usage and all three categories of post-task perceptions of the LLM: Confidence (), Accuracy (), and Helpfulness ().
+However, there is significant positive association between frequency of usage and all three categories of post-task perception: confidence (), accuracy (), and helpfulness ().
 This suggests that while people who were more familiar and open to using LLM tools generally hold higher view of ChatGPT’s advice (across all conditions), it did not impact actual rates of reliance in the task.